### v1

# Quantum mechanics provides the physical basis of teleological evolutions

Abstract
We show that the quantum computational speedup of quantum algorithms is due to their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Naturally, this is a destabilizing result: the teleological notion has long been rejected by science and even, for its finalistic character, viewed as a belief. We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolutions of the living for which the teleological notion was conceived. This work should show that the discovery of the quantum computational speedup confers full scientific dignity to the holistic and finalistic view of physical reality traditionally opposed by science.
1 Introduction
According to natural philosophy, teleological evolutions are evolutions attracted by a final cause; obviously, it would be causation from the future, i.e. retrocausation. Before the advent of modern science, the teleological notion was applied to the evolutions of life, where it provided a potentially enlightening relation between evolution and functional structure of the living: teeth were created for the purpose of chewing. As is well known, with the advent of modern science the teleological notion was dismissed for the alleged absence of a physical basis; for its finalistic character, it was also ascribed to beliefs.
In this work, we shall show that the relatively recent discovery of the quantum computational speedup [^1] By quantum (computational) speedup, we will always mean its certain form, the fact that a quantum algorithm certainly solves its problem with fewer logical operations (computation steps) than possible in the classical world. should change this position of science. This is because quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again). Since this will turn out to be surprisingly evident, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to a deeply rooted rejection, on the part of physicists, of the related notions of retrocausality and finalism. The notion of teleological evolution, of course, involves both.
We, with coauthors, already went against this possible refusal in the evolutionary approach to quantum speedup [1, 2, 3, 4, 5, 6]. In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obey the following rule:
In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
We called the rule in question the advanced knowledge rule. Note that knowing in advance one of the possible halves of the information about the solution one will produce and measure in the future and benefiting from this knowledge to produce the solution with fewer computation steps is a causal loop. We can therefore say that, according to this rule, the quantum speedup is explained by a mutually exclusive or of causal loops, one for each possible way of taking half of the information about the solution.
We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question. Also the notion of finalism acquires a precise definition. The evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem); let us underline the novelty of this concept. Under the advanced knowledge rule, in addition to responding to the philosophical notion of teleological evolution, quantum algorithms obviously provide a scientific example of it.
Until now, the advanced knowledge rule was, so to speak, an empirical rule satisfied by all the major quantum algorithms. In the present work, we should have demonstrated it, incidentally, in a surprisingly simple way. Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
We now go to the evolutions of living to which the teleological notion was first applied. Naturally, one could see an enormous difference between quantum algorithms and the evolutions of living. The former have maximum durations of a few dozen microseconds and a rather limited number of quantum bits, on the order of one thousand; beyond that, today, there is quantum decoherence. The latter have a duration of billions of years and can have the size of the Earth. However, interestingly, the present context suggests a way out. The difference is enormous in the usual classical way of viewing macroscopic reality but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way. Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such (as we shall see), it can occur with quantum speedup and thus be teleological in character.
To see that the evolutions of living have the form of quantum algorithms, in the first place, we should extend the usual quantum description of quantum algorithms, which is limited to the unitary computation of the measurement of the solution, to the upstream process of setting the problem. We also need some terminology. Consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers; we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. Of course, the problem-setting (the number of the drawer with the ball) is here also the solution of the problem. The extended quantum description of the quantum algorithms is thus as follows: an initial measurement randomly selects the problem-setting out of a quantum superposition (or, indifferently, mixture) of all the possible problem settings, the unitary computation of the corresponding solution builds up one-to-one correlation between the problem-setting selected by the initial measurement and the solution, the final measurement reads the solution.
Note that, according to the advanced knowledge rule, in the optimal quantum case, the number of computation steps needed by the quantum algorithm is that needed in the optimal logical case benefiting from the advanced knowledge of half of the information about the solution. Also note that, however, this is just the room for the quantum speedup. The quantum algorithm can fully benefit of it, or partly, or not at all.
Now we go to the evolution of the universe under the fundamental physical laws and the fundamental constants. The version of the Anthropic Principle we are dealing with is about the fact that the least change of the value (for values) that the fundamental physical constants have in the universe we are in would have produced a relatively trivial universe unable to develop life. This can be formulated as follows.
We should think of a plurality of universes, all with the same fundamental laws and each with its own value of the fundamental constants. We consider any one of these universes, in the mutually exclusive or of all the possible universes.
The Fine-tuned Universe version of the Anthropic Principle says that its evolution under the same fundamental physical laws and its own value of the physical constants, builds up a (say) one-to-one correlation between the value of the fundamental constants since the beginning of time and the “value of life” (in particular, “existence of sentient life” and “absence of life”).
Under the quantum cosmological assumption, the evolution of the universe toward the value of life becomes unitary quantum and the one-to-one correlation becomes quantum. Under the quantum cosmological reading, the plurality of universes becomes their quantum superposition and the selection of any one universe in the mutually exclusive or of all universes can be performed by an initial measurement (see what follows) of the value of the fundamental constants in the quantum superposition of all the possible universes.
This is formally identical to a quantum algorithm (see the above extended quantum description of it). Therefore, as we shall see, it can occur with quantum speedup, in the sense that there is certainly the room for its occurring with quantum speedup. Correspondingly, it can be teleological in character.
By the way, the notion of measurement requires that of an observer, and, no matter the value of the fundamental constants, at the beginning of time there could have been none. However, we can do without this initial measurement. There would then be a unitary evolution of the quantum superposition of all possible universes. In the universe we are in, and at the time of today, there would be the sentient observer who measures the value of life also selecting the corresponding value of the fundamental constants. According to the version of the Anthropic Principle we are dealing with, at the time of today these two observables would already be maximally entangled. Since the value of the fundamental constants never changes along the unitary evolution of the universe, all is as if it were selected at the beginning of time.
Summarizing, in the universe we are in, its evolution toward life can have occurred with quantum speedup (in the sense that there is certainly the room for this) and thus have been teleological in character. We shall also see that the same would consequently hold for the evolution of life on Earth (in principle, the trace over all the rest of that of the universe toward life) and the evolutions of species.
By the way, all the above, the quantum cosmological assumption and the consequent reading of the Fine-tuned Universe version of the Anthropic Principle, can, of course, be considered arbitrary. We would argue that it is anyway a plausible possibility, thus very different from the current notion that a teleological evolution of life is an absurdity.
Of course, if taken seriously, the above would also be highly destabilizing. Besides providing a unified quantitative explanation of the quantum speedups, the present result would change the very foundations of the natural sciences, restoring what could be an enlightening relation between evolution and functional structure of the living. Again, our purpose here is showing that quantum algorithms are certainly teleological evolution, and their distance from the evolutions of life could even be zeroed under a plausible way of viewing macroscopic reality, i.e. the quantum cosmological assumption.
Also note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science. In the present theory of the quantum computational speedup, they could acquire a physical basis.
2 Time-symmetric explanation of the quantum speedup
The present way of thinking about quantum algorithms and their quantum speedups has radical differences from the mainstream one. In this section, we provide it at a conceptual level, but always making a point-by-point reference to the formalizations, all postponed in the Appendix. The section is divided into subsections corresponding to the various arguments.
Let us first acknowledge the fact that the discovery of the quantum speedup has been one of the fundamental discoveries of the past century, which, together with quantum communication, has created an entirely new branch of physics. However, having said this, for the sake of clarity, we will be completely frank in criticizing what we consider to be the shortcomings of the current way of viewing quantum computation.
2.1 On the explanation of the quantum speedup
Let us begin with the advanced knowledge rule, which we repeat here for convenience.
In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
Of course, this rule offers a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information. Let us discuss this point. To the best of our knowledge, the absence of such an explanation has never been acknowledged by the scientific community.
People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition. Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, cannot be considered its explanation. By the way, what distinguishes science from opinion is, in particular, its quantitative character. The fact that the most popular explanation of the quantum speedup, a quantitative feature, is qualitative in character, would seem to be rather unscientific. The only justification we can see is that this is a way of removing the fact that, forty years after its discovery—an explanation of the quantum speedup is still missing. This is naturally our opinion. However, as a matter of fact, outside our (destabilizing) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
Neither can be considered an explanation the fact that each quantum algorithm naturally explains its quantum speedup. In fact, it would be an explanation completely different for different quantum algorithms. It would be like having different explanations for different throws of a stone without knowing that they all describe a parabola. Obviously, unification is a fundamental objective of any description of physical phenomena.
In our view, for what concerns the current mainstream explanation of the quantum speedup, this is the case of crying out: the emperor is naked! Our emphasis on this is because it could balance the high unconventionality of our proposal.
Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation. However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against a deeply rooted de facto ban on retrocausality. In this work, we should have demonstrated it, indeed in a surprisingly simple way. As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the destabilizing character of the rule.
2.2 First demonstration of the advanced knoledge rule
In any case, in order to demonstrate the advanced knowledge rule, we should first amend the current quantum description of quantum algorithms: as it is, it would completely obscure both demonstrations. The quantum description in question is limited to the unitary computation of the solution, where the quantum speedup resides, and the measurement of the latter. Of course, a canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement. The current quantum description is obviously focused on the central problem of how to achieve a quantum speedup with respect to to what is logically possible. By the way, the explanation of the quantum speedup we will be dealing with completely ignores the way in which it is reached, it is limited to the fact that it is reached.
Anyway, the above shortcoming is simply remedied: we should add to the current quantum description of the process of solving the problem that of the upstream process of setting the problem. An initial measurement in the quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random [^2] The problem-setter could unitarily change it at will; this, here, would change nothing., the unitary computation and the final measurement of the corresponding solution follow—for the formalization of this, see Section (1) of the Appendix. We would recommend to look at it; it is trivial but introduces all the definitions that will be used in the subsequent formalizations.
The quantum description of the quantum algorithm defined until now is with respect to the customary observer external to the quantum process. To present ends, we should relativize it with respect to the problem-solver, who computes the solution of the problem. Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the quantum description of the quantum algorithm with respect to her. From her, the problem-setting selected by the initial measurement should be concealed. As we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action. This concealment is obtained by postponing to the end of her unitary computation of the solution the projection of the quantum state associated with the initial measurement; see Section (2) of the Appendix.
Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”. It is ordinary because it is still the customary “causal” description of mainstream quantum mechanics—where causality always goes in the same time-direction.
The first demonstration of the advanced knowledge rule simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
By the way, involving only the measurements of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible. Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known. Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [7]. Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup. (ii) In Long’s [20] version, it always provides the solution with certainty; this simplifies things. (iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization. (iv) It is optimal in character; this, of course, constitutes an important reference.
We show how a classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule. For the moment, we consider the simplest instance of Grover algorithm. As we shall soon see, everything immediately generalizes to the full fledged Grover algorithm and most general forms of quantum correlation (in quantum algorithms, between the problem-setting and the solution of the problem).
The problem to be solved is as follows. Bob, the problem-setter, hides a ball in a chest of four drawers. Alice, the problem-solver, is to locate the ball by opening drawers. In the worst classical case, she has to open three drawers. If the ball is only in the third drawer opened, she has solved the problem. If not, it must be in the only drawer not yet opened, and she has solved the problem of locating the ball as well.
Instead, Grover algorithm (its simplest instance) always solves the problem by opening just one drawer, in a quantum superposition of the four possible drawer numbers. Obviously, always solving the problem by opening just one drawer would be impossible in the classical case: indeed, we are dealing with a quantum computational speedup.
There is a classical logic equivalent to the fact that the quantum algorithm always solves the problem by opening a single drawer. By the way, since the fact that a single drawer is opened is logically true in itself, we can forget the specification that this is done for a quantum superposition of the four drawer numbers.
The logical equivalence is as follows. Logically, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers (one of which, of course, the one with the ball) and took advantage of this knowledge to solve the problem by opening just one of them. In fact, if the ball is in the opened drawer, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. This also means knowing in advance half of the information that specifies the solution of the problem that will be produced and measured in the future.
Let us open a parentheses to observe that we are not talking about the quantum algorithm as it were a classical algorithm. In other words, here, classical logic should not be confused with classical physics, which might be a temptation in the present case.
Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer. Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, for her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that the information in question retrocausally comes to her from her future measurement of the solution. This is formalized in Section (3) of the Appendix. However, before looking at it, one should read the description of the time-symmetrization of the quantum algorithm that we shall provide in a moment.
Note that we are dealing with a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings). There is a causal loop for each way of taking half of the information about the solution. Therefore, we can say that this quantum speedup is logically explained by the mutually exclusive or of the causal loops we are dealing with. Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup. In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.
This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers. As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup. This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting of this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers. Of course, this is what the optimal Grover algorithm does. Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm.
Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers. In still equivalent terms, the dimension of the Hilbert in which the search for the solution occurs reduces from $N$ to $\sqrt{N}$.
Holding for Grover algorithm, the above retrocausal explanation of the quantum speedup must hold for the process of solving with a quadratic quantum speedup any NP problem. Here we do not care about the fact that most NP problems are classically solved with a number of computation steps exponential in their size, so that a quadratic quantum speedup can be seen as a little advantage. Of course, for the present ends, we are just interested in having a quantum speedup with respect to what is classically possible.
We shall see further generalizations in the following of this section. In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.
2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete. Of course, to bring it in line with the classical logic description, it should be completed. It turns out that this is simply done by time-symmetrizing it.
We describe this operation. For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm. We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of a quantum algorithm, all the information is selected by the initial measurement. Naturally, there is a plurality of possible ways of evenly sharing: we should take their quantum superposition.
In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement. That performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement. As we can see in Section (3) of the Appendix, in this new initial measurement outcome it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps. This is the physical correspondent of one of the logical causal loops of the advanced knowledge rule.
Eventually, we should take the quantum superposition of all the time-symmetrization instances. The time-symmetrized quantum description of the quantum algorithm (i.e. its completed quantum description) becomes the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the quantum speedup: So completed, also the quantum description of the quantum algorithm describes the advanced knowledge rule described by its classical logic description—it describes its physical counterpart.
2.4 Second demonstration of the advanced knowledge rule
The second demonstration of the advanced knowledge rule simply comes from the observation that the time-symmetrization of the ordinary quantum description of the quantum algorithm, which provides the (physical counterpart of the) advanced knowledge rule, is mandatory also for another reason, as follows.
In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting. Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it is simultaneous with the final measurement of the solution. In this case, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements. Of course, there would be no reason why a measurement selected more information than the other. As we have just seen, this time symmetrization generates the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule. This is indeed the second demonstration of the advanced knowledge rule.
By the way, this result is reminiscent of Wheeler and Feynman’s [8] time-symmetric absorber theory; we would say: with the advantage that the time-symmetrization we are presently dealing with relies on a reversibility of the quantum process that, unlike the classical one, is such in both an ideal and objective way.
2.5 Origin of the retrocausality
A quantum algorithm is a relatively complex process. A natural question is whether the retrocausality implicit in the classical logic description of its overall result, i.e. the quantum speedup, derives from some more elementary part of it. Of course, in its ordinary quantum description, the unitary computation of the solution is an alternation of generations of quantum state superpositions and quantum interferences. Since the latter have a classical analog, the essentially quantum generation of a quantum state superposition is the suspect.
Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality. We resort to a prescientific notion of state superposition. It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this. Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities. Until now, being on the two banks at the same time would seem to be a logical impossibility. However, let us note that it would be impossible in a classical logic description closed to retrocausality. Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^3] By the way, classical logic is neutral for what concerns the notion of causality. Customarily associating it with a causality that always goes in the same time-direction must be consequent to the de facto no-retrocausality precept. In principle, classical logic should also go along with the coexistence of causality and retrocausality.. This said, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path. This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
Let us now bring the above classical logic description of a state superposition to a quantum state superposition. Think of a photon going in quantum superposition through the two branches of a Mach-Zehnder interferometer. The two branches should be tuned in such a way that the photon always ends up in the same detector, so that we have a (of course, time-reversible) unitary evolution that goes from before the photon enters the interferometer to immediately before its detection.
The classical logic description of this can be as follows. In entering the interferometer, the photon takes, say, the upper branch until it reaches the state immediately before detection (by the detector in question). Of course, this also provides the phase of the photon along the path followed. Then (in causal order), the photon goes backward in time undoing the path just followed until before entering the interferometer. Then, it goes forward in time again taking the lower branch until reaching the detector in question and, this time, being detected. The zigzag path followed in this way by the photon through the interferometer, with the associated photon phases, of course, describes in a classical logic way the fact that the photon goes through the interferometer in a quantum superposition of the upper and lower branches, eventually ending up in that detector. It should be clear that, also in this case, the involvement of retrocausality in the classical logic description is necessary.
We should think that the classical logic description of the sequence of generations of quantum state superpositions and interferences that constitutes the unitary computation of the solution, in the end, simplifies into a description that implies the advanced knowledge rule.
Summarizing, at the root of the retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it. By the way, note that this is not an idle speculation, as might even seem. Indeed, we are dealing with a retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
2.6 Generalization
In the second demonstration of the advanced knowledge rule, we have shown that a unitary evolution between two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it. Note that we never needed to look inside that unitary evolution $\hat{\mathbf{U}}_{1,2}$: see Tables I through IV of the Appendix. Therefore, this time-symmetrization does not only hold for quantum algorithms, it lends itself to an important generalization:
The ordinary quantum description of any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it.
Let us call it the quantum correlation time-symmetrization rule.
The case where the one-to-one correlation is built up by the unitary evolution characterizes the quantum search algorithm, where this building up occurs with quadratic quantum speedup. However, the possibility of occurring with quadratic quantum speedup holds for any unitary evolution that responds to the rule in question and builds up one-to-one quantum correlation between two initially unrelated observables. Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says. Of course, in the case that we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality. Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables. With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule. The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops which locally explain it and play the role of the hidden variables envisaged by Einstein et al [9]. Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement—see [5]. Note that this is what Einstein et al required from their hidden variables. That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process.
In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it. Of course, it also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
This ends the first part of the work. The second is about the consequences of the advanced knowledge rule or the quantum correlation time-symmetrization rule. These consequences are both widespread and destabilizing: indeed, the time-symmetric form of retrocausality implicit in these rules yields a different way of viewing physical reality.
The consequence just seen, demonstrating the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [9], is destabilizing for the following reason. After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [10] and Ekert’s quantum cryptography based on it [11], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
Another destabilizing consequence of the advanced knowledge rule is that, according to it, quantum algorithms are teleological evolutions, whereas, currently, science claims that teleological evolutions are without physical basis.
An even more destabilizing consequence is that the physical basis of the teleological character of quantum algorithms, under the quantum cosmological assumption, applies as well to the evolutions of the living. Of course, the teleological notion was once applied by natural philosophy to the evolution of life; by natural philosophy, it is also applied today [12]. As already mentioned, with the advent of modern science, it was dismissed for the alleged absence of a physical basis. For its finalistic character, it has also been ascribed to the category of beliefs. The next subsection is dedicated to the teleological notion under the present theory of the quantum computational speedup.
2.7 The physical basis of teleological evolutions
Let us repeat another time the advanced knowledge rule, which naturally implies the teleological character of quantum algorithms.
In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps. The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
We have demonstrated this rule in two redundant ways. The first, by describing the quantum speedup in a classical logic way, the second by a time-symmetrization of the ordinary quantum description of the quantum algorithm which is mandatory to satisfy the atemporal symmetry implicit in it: we have seen that this time-symmetrization generates the physical correspondent of the advanced knowledge rule. With this, of course, we have also demonstrated the teleological character of quantum algorithms: indeed, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
We have also seen that, under the quantum cosmological assumption that the universe evolves in a unitary quantum way, and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the value of life acquires the form of a quantum search algorithm. As such, it can occur with a quadratic quantum speedup. The “can occur” is in the sense that there is certainly room for a quantum speedup, but the unitary evolution must benefit of it; it should occur with a just suboptimal quantum speedup if the search for the value of life is random [21] .
We go now to the fact that this possible teleological character would be inherited by the evolution of life on Earth and the evolution of species.
Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life applies to the evolution of the universe toward life. Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth. In principle, the latter evolution is derivable from the former evolution by tracing out all the rest. Therefore, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life. Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe. Second, the amount of physical resources used by the evolution of life on Earth cannot exceed the amount of physical resources used by its image in the evolution of the universe. In fact, the latter resources comprise the former. In summary, the evolution of life on Earth could occur with quantum speedup (in the sense that there is certainly the room for that) and consequently be teleological in character; it should actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random. Then, just for consistency, the evolutions of species would also be the same.
Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would mean for the notion of the evolutions of species. It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs that did not fly. Very light, strong but flexible, cave and air-proof, even at that time feathers were ideal for flight, very close to those of the future birds. So why anticipate something designed for flight many millions of years in advance?
The usual explanation, of course, is causal. First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually, feathers turned out to be what was needed to conquer the air.
All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred. See our preprint [6] for a more comprehensive discussion of the teleological notion.
All this is obviously destabilizing, it is a way of viewing the macroscopic reality that goes against the modern way of viewing it, and revamps ancient natural philosophy views which were ascribed by science to completely unscientific beliefs.
Let us then discuss more extensively this contrast between the scientific and the natural philosophy view of the teleological notion.
The notion of teleology as final cause was introduced almost two and a half millennia ago by Plato and Aristotle to explain the “formation” of the functional structure of the living (Aristotle: teeth were created for the purpose of chewing). Of course, at the time, there was not yet the notion of evolution of the living. In more recent times, the teleological notion was applied by natural philosophy to the evolutions of the living—by natural philosophy it is also applied today [12]. It was then dismissed by science for the alleged absence of a physical basis—precisely because it implies a retrocausality deemed to be physically impossible. Of course, at the time of classical physics, this was completely justified. For its finalistic character, the teleological notion was also ascribed to the category of beliefs.
It should however be noted that this dismissal was not without regret; indeed, it meant doing without a possibly enlightening relation between the notion of teleological evolution of life and that of the functional structure of living (teeth were created for the purpose of chewing). We cite two authoritative objections to this dismissal. One is a famous sentence by the evolutionary biologist Haldane [26]:
Teleology is like mistress to a biologist, he cannot live without her, but is unwilling to be seen with her in public.
The other is the position of the natural philosopher Nagel [12], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view. In our possibly biased opinion, this would be a case in which common sense prevails over a de facto no-retrocausality precept that, according to the present work, would be a prejudice.
This also raises the question of how to describe a macroscopic evolution endowed with quantum speedup. Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
We quote a precedent in the quantum description of macroscopic natural evolutions. It is about the asymptotic behavior of predator species competing for the same prey [27]. It is explained by a quantum model, whereas it would be difficult to explain it in a classical way. In our case, the canonical way of describing the evolution of life on Earth would be to obtain it from the description of the evolution of the universe toward life by tracing out all the rest. Since this obviously holds only in principle, we should also model the macroscopic teleological evolutions of species endowed with quantum speedup in a quantum way. In other words, we should accept quantum superposition, entanglement, and correlation between the observables of macroscopic objects.
This might even be a good approximation, since, in the universe we are in, the evolution of life on Earth in its time lapse should have been almost autonomous from the rest of the unitary quantum evolution of the universe toward life. Indeed, if it were completely autonomous, of course, it would be a unitary quantum evolution itself.
Incidentally, it might also be interesting to derive the description of the evolution of life on Earth from that of the universe toward life by tracing out all the rest at the level of toy models of the two evolutions. Within the present theory, this might highlight significant characters of a possible teleological evolution of life on Earth.
By the way, of course, the quantum cosmological view of the evolution of the universe would “unify” quantum and classical physics simply by canceling the latter. However, naturally, classical physics turns out to work extremely well in describing macroscopic reality. Presently, it would have the shortcoming of being unable to describe macroscopic teleological evolutions. This situation might point to a particular object of research. It could be interesting to study how the operation of “tracing-over the rest” (at the toy model level) reconciles a teleological behavior in the long term with a classical behavior in the short term.
Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup. According to the present theory, it would be an even more fundamental discovery than currently thought. The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [15] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—an attempt to unify relativity and quantum theory. Then, there was the introduction of the notion of logically reversible computation by Bennett [16] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions. The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [17]. Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [18]: with it, we have the logically reversible gates needed to perform quantum computations. The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [19] in 1985.
Note how the notion of the logical reversibility of computation played an essential role in the development of the notion of quantum computation. The present time-symmetric quantum description of quantum algorithms could be a further step in this direction. Involving a time-symmetric coexistence of the two possible time-directions of causality, it could be said that it exploits the objectively ideal time-reversibility of unitary quantum evolutions in the possibly deepest way.
3 Precedents
We have seen that applying the quantum correlation time-symmetrization rule to quantum nonlocality provides a local explanation of it. Now, we recall a historical precedent of this.
It is a work by Costa de Beauregard [22] that also provides a retrocausal explanation of quantum nonlocality. In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement. Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar. Interestingly, at the time, Costa de Beauregard was working for de Broglie, who forbid him to publish that work involving retrocausality for years. Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory. See [23] for an account of this episode. It is rather significant of the attitude, of even top physicists, toward the retrocausality notion.
Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [24, 25]. Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics. However, first, none of them concerns the quantum computational speedup and, second, they all apply to very specific physical situations to which they might be confined. On the contrary, the present time-symmetric form of retrocausality should be sort of universal for quantum correlations. The possible relation of those works to the present work, as things are now, would not be evident. Of course, it could be the object of further investigation.
4 Conclusions
The present work goes against a long-standing de facto no-retrocausality tradition of physics. More specifically, since causality could well go backward in time along a unitary quantum evolution that is both objectively and ideally time-reversible, we should say that, in the quantum world, the scientific tradition amounts to asserting the mutual exclusivity between the two possible time directions of causality.
Let us first address the reliability of the work, a key issue because of its unconventionality.
That the classical logic description of the quantum speedup must resort to retrocausality should be obvious. Indeed, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers. To the problem-solver, the only possible source of this information is the solution of the problem she will produce and measure in the future: hence retrocausality. We have seen how this immediately generalizes to the full fledged Grover algorithm and a most general form of quantum correlation.
At this point, we should address the legitimacy of applying classical logic to what the quantum algorithm does. Let us recall our two arguments in favor of this choice.
The first is that, only involving the measurements of commuting observables (the problem-setting and the solution of the problem), quantum algorithms (and what they do) should be subject to classical logic.
The other is that all the results obtained under the assumption that quantum algorithms can be described by classical logic are also obtained by time-symmetrizing their ordinary quantum description; this, in turn, is mandatory to satisfy the atemporal symmetry between the simultaneous measurements of two maximally entangled observables implicit in the ordinary quantum description of quantum algorithms.
We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences (or right for them).
The other side of the coin of the destabilizing character of the work is, of course, its revolutionary character, always due to the retrocausality implicit in it. We should distinguish between what should be certain and what could be just plausible. Beginning with what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedup, which, forty years after its discovery, was still missing from mainstream quantum mechanics. (ii) It theoretically confirms the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it. (iii) It highlights the teleological character of quantum algorithms; by the way, given the negative view that the teleological notion has in science today, and the scope of this notion, this should be the most revolutionary result of all. About the plausible things: It shows that, under a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could apply as well to the evolutions of the living.
Let us eventually note that all this should change our way of seeing physical reality. Of course, for this, it should be sufficient the revamping of the teleological notion. Why now? Naturally, after the relatively recent discovery of the quantum computational speedup.
Appendix
We have seen that, according to the present theory, the usual quantum description of quantum algorithms is incomplete. How to complete it has already been discussed at a conceptual level in the previous part of the work. We now provide the formalization. We do it in the simplest instance of Grover algorithm; we shall see that everything immediately generalizes.
We should perform the following steps:
(1) Extend the usual quantum description of the quantum algorithm, limited to the process of solving the problem, to the upstream process of setting the problem: One should measure the problem-setting in a quantum superposition of all the possible problem-settings. The unitary computation and measurement of the corresponding solution follow.
(2) Relativize the extended quantum description of the quantum algorithm with respect to the problem-solver, from whom the problem-setting selected by the initial measurement should be concealed (it would tell her also the solution of the problem, before she begins her problem-solving action). Note that the advanced knowledge of one of the possible halves of the information about the solution of the problem must be in the quantum description of the quantum algorithm with respect to the problem-solver.
(3) Complete the quantum description of the quantum algorithm obtained so far by time-symmetrizing it.
Steps (1) through (3) are developed in the following correspondingly numbered sections.
(1) Extending the usual quantum description of quantum algorithms to the process of setting the problem
The usual quantum description of quantum algorithms is limited to the unitary process of computing the solution of the problem and the final measurement of the solution. We extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition of all the possible problem settings.
We introduce the notation; we make reference to the simplest, four drawer, instance of Grover algorithm:
We number the four drawers in binary notation: $00,01,10,11$.
We will need two quantum registers of two quantum bits each:
A register $B$ under the control of the problem-setter Bob contains the problem-setting observable $\hat{B}$ (the number of the drawer in which the ball is hidden) of eigenstates and corresponding eigenvalues:
$|00\rangle_{B}$ and $00$; $|01\rangle_{B}$ and $01$; $|10\rangle_{B}$ and $10$; $|11\rangle_{B}$ and $11$.
Measuring $\hat{B}$ in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random.
A register $A$ under the control of the problem-solver Alice contains the number of the drawer opened by her. The content of register $A$ is the observable $\hat{A}$ of eigenstates and corresponding eigenvalues:
$|00\rangle_{A}$ and $00$; $|01\rangle_{A}$ and $01$; $|10\rangle_{A}$ and $10$; $|11\rangle_{A}$ and $11$.
Measuring $\hat{A}$ at the end of the quantum algorithm yields the number of the drawer with the ball selected by the initial measurement, namely the solution of the problem. Note that $\hat{B}$ and $\hat{A}$ commute.
The quantum description of the extended quantum algorithm is:
$\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&&\\ \Downarrow&&\\ |01\rangle_{B}\left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table I)
Here and in the following we disregard normalization.
We should start from the quantum state in the upper-left corner of Table I diagram and then follow the vertical and horizontal arrows. Initially, at time $t_{1}$, both registers, each by itself, are in a quantum superposition of all their basis vectors. By measuring $\hat{B}$ (the content of $B$) in this initial state, Bob selects the number of the drawer with the ball at random; say it comes out $01$: see the state of register $B$ under the vertical arrow [^4] By the way, Bob would be free to unitarily change at will the sorted out number of the drawer with the ball. We omit this operation that would change nothing here. .
It is now Alice’s turn. The unitary part of her problem-solving action—by the unitary transformation $\hat{\mathbf{U}}_{1,2}$—produces the solution of the problem, of course, never changing the problem-setting.
Note that we will never need to know $\hat{\mathbf{U}}_{1,2}$ (i.e. Grover algorithm). We will only need to know that, between the outcome of the initial measurement and that of the final measurement, there is a unitary evolution that builds up one-to-one correlation between the two initially unrelated observables $\hat{B}$ and $\hat{A}$.
At the end of Alice’s problem-solving action, at time $t_{2}$, register $A$ contains the solution of the problem, i.e. the number of the drawer with the ball—bottom-right corner of the diagram.
Eventually, by measuring $\hat{A}$ (the content of register $A$), which is already in one of its eigenstates, Alice acquires the solution with probability 11—the quantum state in the bottom-right corner of the diagram naturally remains unaltered throughout this measurement.
(2) Relativizing the extended quantum description of the quantum algorithm with respect to the problem-solver
We must relativize the quantum description of the quantum algorithm of Table I, which is with respect to the customary external observer, with respect to the problem-solver. This amounts to concealing from her the projection of the quantum state associated with the initial measurement of the problem-setting
Since Alice’s problem-solving action never changes the number of the drawer with the ball selected by the initial measurement and the observables $\hat{B}$ and $\hat{A}$ commute, describing this concealment is simple. It suffices to postpone after the end of Alice’s problem-solving action the projection of the quantum state associated with the initial measurement of $\hat{B}$ (or, indifferently, to postpone the very measurement of $\hat{B}$). The result is the following quantum description of the quantum algorithm with respect to Alice:
$\begin{array}[]{ccc}\begin{array}[]{c}\text{time}\text{ }t_{1}\text{, meas. of}\;\hat{B}\end{array}&t_{1}\rightarrow t_{2}&\text{\text{time }\text{}}t_{2}\text{, meas. of}\;\hat{A}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ &&|01\rangle_{B}|01\rangle_{A}\end{array}$ (Table II)
To Alice, the initial measurement of $\hat{B}$ leaves the initial state of register $B$ unaltered. At the beginning of her problem-solving action she is completely ignorant of the number of the drawer with the ball selected by Bob—see the state of register $B$ on the left of $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$.
The unitary transformation $\hat{\mathbf{U}}_{1,2}$ is then performed for a quantum superposition of the four possible numbers of the drawer with the ball. The state at the end of it—on the right of the horizontal arrows—is a quantum superposition of tensor products, each a possible number of the drawer with the ball in register $B$ tensor product the corresponding solution (that same number) in register $A$. Since $\hat{B}$ and $\hat{A}$ commute, the final measurement of $\hat{A}$ must project this superposition on the tensor product of the number of the drawer with the ball initially selected by Bob and the corresponding solution—bottom-right corner of Table II diagram.
Note that this relativization relies on relational quantum mechanics [28, 29] where the quantum state is relative to the observer. In the following, we will call this quantum description of the quantum algorithm with respect to the problem-solver its ordinary quantum description.
(3) Time-symmetrizing the ordinary quantum description of the quantum algorithm
As anticipated in the Introduction, the ordinary quantum description of Grover algorithm is incomplete and is completed by time-symmetrizing it. We should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes evenly shares between the initial and final measurements. The half information selected by the initial measurement should propagate forward in time, by $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement. Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time, by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$, until replacing the previous outcome of the initial measurement.
We consider the time-symmetrization instance where the initial measurement of $\hat{B}$ reduces to the measurement of $\hat{B}_{l}$, the left bit of the two-bit number contained in register $B$; say that it randomly selects left bit 0. The final measurement of $\hat{A}$ should correspondingly reduce to that of $\hat{A_{r}}$, the right bit of the two-bit number contained in register $A$; say that it randomly selects right bit $1$. The randomly selected number of the drawer with the ball is thus $01$.
In this particular time-symmetrization instance, the ordinary quantum description of the quantum algorithm of Table II becomes:
$\begin{array}[]{ccc}\begin{array}[]{c}\;\text{time }t_{1}\text{, meas. of}\;\hat{B}_{l}\end{array}&t_{1\rightleftarrows}t_{2}&\text{time }t_{2}\text{,meas. of}\;\hat{A_{r}}\\ &&\\ \begin{array}[]{c}\left(|00\rangle_{B}+|01\rangle_{B}+|10\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&\begin{array}[]{c}|00\rangle_{B}|00\rangle_{A}+|01\rangle_{B}|01\rangle_{A}+\\ |10\rangle_{B}|10\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}\\ &&\Downarrow\\ \begin{array}[]{c}\left(|01\rangle_{B}+|11\rangle_{B}\right)\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table III)
The projection of the initial state of register $B$ due to the measurement of $\hat{B}_{l}$—top-left corner of the diagram—is postponed after Alice’s problem-solving action, namely outside the present table which is limited to that action. With this postponement, the initial state of register $B$ goes unaltered through the initial measurement of $\hat{B}_{l}$ becoming, by $\hat{\mathbf{U}}_{1,2}$, the quantum superposition of tensor products above the vertical arrow. Alice’s final measurement of $\hat{A_{r}}$, which in this time-symmetrization instance selects the value $1$ of the right bit of register $A$, projects the state above the vertical arrow on the state below it. Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement. This backward in time propagation is an instance of the time-symmetrized description of the quantum algorithm with respect to the problem-solver. As we can see from the diagram, it tells her, before she begins her problem-solving action, one of the possible halves of the information about the solution, namely that the ball is in drawer either 01 or 11: look at the state of register $B$ in the bottom left corner of the table.
In summary, the causal loop is the zigzag diagram of Table III. We can see that it is one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. The state of register $B$ in the top-left part of the diagram tells us that Alice, before beginning her problem-solving action, is completely ignorant of the number of the drawer with the ball. Through the zig-zag, this state changes into the state of register $B$ in the bottom-left part of the diagram, which tells us that Alice, before beginning her problem-solving action, knows in advance that the ball is in drawer either $01$ or $11$; we are dealing with an effective causal loop that changes its own past. She can then solve the problem in an optimal logical way by opening just one of the two drawers, indeed as in the classical logic description of what the quantum algorithms does: look at the bottom line of the diagram in Table III.
Note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, this bottom line of Table III diagram (mathematically) can also be read in the usual left to right way:
$\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
We can see that, in each time-symmetrization instance (causal loop) the computational complexity of the problem to be solved by Alice quadratically reduces from the problem of locating the ball in one of $4$ drawers to that of locating it in $2=\sqrt{4}$ drawers. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces from $4$ basis vectors to $2=\sqrt{4}$ basis vectors.
Eventually, we should take the quantum superposition of all the time-symmetrization instances.
Note that this time-symmetrization, mathematically, leaves the unitary computation of the solution, $\hat{\mathbf{U}}_{1,2}$, unaltered, but changes the behavior of causality along it. The single causal loop of Table II is replaced by the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup.
In subsection 2.6, we have seen that all this, first, immediately extends to any number of drawers, then, to more general forms of quantum correlation.
Acknowledgements
Thanks are due to Eliahu Cohen, Artur Ekert, Avshalom Elitzur and David Finkelstein for useful comments, Daniel Sheehan for organizing the San Diego AAAS Pacific Division conferences on retrocausality, a far-looking forum for the discussion of frontier, also unorthodox, physics, and Mario Rasetti for organizing the first series of international workshops on quantum communication and computation, the 1991–96 Elsag Bailey–ISI Turin Villa Gualino workshops.

### v2

@@
 Abstract
-We show that the quantum computational speedup of quantum algorithms is due to their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
-Naturally, this is a destabilizing result: the teleological notion has long been rejected by science and even, for its finalistic character, viewed as a belief.
-We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolutions of the living for which the teleological notion was conceived.
-This work should show that the discovery of the quantum computational speedup confers full scientific dignity to the holistic and finalistic view of physical reality traditionally opposed by science.
+We show that the quantum computational speedup is due to the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
+We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living for which the teleological notion was originally conceived.
 1 Introduction
-According to natural philosophy, teleological evolutions are evolutions attracted by a final cause; obviously, it would be causation from the future, i.e.
-retrocausation.
-Before the advent of modern science, the teleological notion was applied to the evolutions of life, where it provided a potentially enlightening relation between evolution and functional structure of the living: teeth were created for the purpose of chewing.
-As is well known, with the advent of modern science the teleological notion was dismissed for the alleged absence of a physical basis; for its finalistic character, it was also ascribed to beliefs.
-In this work, we shall show that the relatively recent discovery of the quantum computational speedup [^1] By quantum (computational) speedup, we will always mean its certain form, the fact that a quantum algorithm certainly solves its problem with fewer logical operations (computation steps) than possible in the classical world.
-should change this position of science.
-This is because quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
-Since this will turn out to be surprisingly evident, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to a deeply rooted rejection, on the part of physicists, of the related notions of retrocausality and finalism.
-The notion of teleological evolution, of course, involves both.
-We, with coauthors, already went against this possible refusal in the evolutionary approach to quantum speedup [1, 2, 3, 4, 5, 6].
-In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obey the following rule:
+Let us briefly examine the very long history of the teleological notion—attraction by a final cause.
+Plato and Aristotle used it to explain the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution.
+With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy.
+It was eventually dismissed by science for the alleged absence of a physical basis.
+All this was before the advent of quantum mechanics.
+Today, the situation of the teleological notion is somewhat complex.
+To the best of our knowledge, outside the present work, there are no physical researches on it.
+Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
+In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science.
+This is because, in the first place, quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
+Since showing this will be surprisingly simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
+We, with coauthors, already went against this position in the evolutionary, retrocausal, approach to the quantum computational speedup [1, 2, 3, 4, 5, 6].
+In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obeyed the following rule:
 In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
@@
 We can therefore say that, according to this rule, the quantum speedup is explained by a mutually exclusive or of causal loops, one for each possible way of taking half of the information about the solution.
+By the way, here and in the following by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible.
+Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
 We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question.
 Also the notion of finalism acquires a precise definition.
-The evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem); let us underline the novelty of this concept.
-Under the advanced knowledge rule, in addition to responding to the philosophical notion of teleological evolution, quantum algorithms obviously provide a scientific example of it.
-Until now, the advanced knowledge rule was, so to speak, an empirical rule satisfied by all the major quantum algorithms.
+The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem).
+Until now, this advanced knowledge rule was, so to speak, an empirical rule shown to be satisfied by all major quantum algorithms.
 In the present work, we should have demonstrated it, incidentally, in a surprisingly simple way.
 Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
-We now go to the evolutions of living to which the teleological notion was first applied.
-Naturally, one could see an enormous difference between quantum algorithms and the evolutions of living.
-The former have maximum durations of a few dozen microseconds and a rather limited number of quantum bits, on the order of one thousand; beyond that, today, there is quantum decoherence.
+As a matter of fact, we have provided two redundant demonstrations of this rule, what should not be despised given its destabilizing character. Let us outline them.
+One is an application of classical logic to what the quantum algorithms does: providing a quantum speedup with respect to what is classically possible.
+By the way, we will have to use a quantum description of the quantum algorithms that is not their customary quantum description.
+The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution.
+It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement.
+We complete the customary quantum description by adding to the description of the process of solving the problem that of the upstream process of setting the problem.
+The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow.
+Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic.
+As we shall see in the next section, the classical logic description of the quantum speedup is indeed the advanced knowledge rule.
+Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete.
+Indeed, it is simply completed by time-symmetrizing it.
+We shall see that the time-symmetrization of the quantum description yields the quantum superposition of the causal loops involved in the advanced knowledge rule.
+In other words, the time-symmetrization of the quantum description of the quantum algorithm generates the physical correspondent of the advanced knowledge rule.
+The second demonstration of the advanced knowledge rule simply comes from the fact that this time-symmetrization of the quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far.
+As we shall see, it is to satisfy an atemporal symmetry implicit in the quantum description.
+Of course, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—it is a demonstration of the rule.
+Of course, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
+A further step, important for what concerns the possible teleological character of the evolutions of the living, is that all the above generalizes to quantum correlation.
+For simplicity, we limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty.
+We can then say that the quantum algorithm builds up one-to-one correlation between two initially unrelated observables (the problem-setting and the solution of the problem) without ever changing the eigenvalue of the observable selected by the initial measurement (the problem-setting).
+We shall show that any building up of quantum correlation of this kind (not only that in quantum algorithms) can occur with quantum speedup and therefore be teleological in character.
+Note that the “can be” means that there is certainly room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation.
+By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
+With this, we can go to the evolutions of the living for which the teleological notion was originally conceived.
+First, let us note that the dismissal of the teleological notion was not without (scientific) regret; indeed, it meant doing without a possibly enlightening relation between the notion of evolution of life and that of the functional structure of the living.
+We cite two authoritative objections to this dismissal.
+One is a famous sentence by the evolutionary biologist Haldane [8]:
+Teleology is like mistress to a biologist, he cannot live without her, but is unwilling to be seen with her in public.
+The other is the position of the natural philosopher Nagel [9], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view.
+In our possibly biased opinion, this would be a case in which common sense prevails over a de facto no-retrocausality precept that, according to the present work, would be a prejudice.
+Let us now go to what the teleological character of quantum algorithms can say to the evolutions of the living.
+Naturally, the first observation must be on the enormous difference between quantum algorithms and the evolutions of the living.
+The former have maximum durations of a few dozen microseconds and a number of quantum bits limited to the order of one thousand; beyond that, in today’s implementations, there is quantum decoherence.
 The latter have a duration of billions of years and can have the size of the Earth.
-However, interestingly, the present context suggests a way out.
-The difference is enormous in the usual classical way of viewing macroscopic reality but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way.
-Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such (as we shall see), it can occur with quantum speedup and thus be teleological in character.
-To see that the evolutions of living have the form of quantum algorithms, in the first place, we should extend the usual quantum description of quantum algorithms, which is limited to the unitary computation of the measurement of the solution, to the upstream process of setting the problem.
-We also need some terminology.
-Consider the following problem.
-Bob, the problem-setter, hides a ball in a chest of four drawers; we call the number of the drawer with the ball the problem-setting.
-Alice, the problem-solver, is to locate the ball by opening drawers.
-Of course, the problem-setting (the number of the drawer with the ball) is here also the solution of the problem.
-The extended quantum description of the quantum algorithms is thus as follows: an initial measurement randomly selects the problem-setting out of a quantum superposition (or, indifferently, mixture) of all the possible problem settings, the unitary computation of the corresponding solution builds up one-to-one correlation between the problem-setting selected by the initial measurement and the solution, the final measurement reads the solution.
-Note that, according to the advanced knowledge rule, in the optimal quantum case, the number of computation steps needed by the quantum algorithm is that needed in the optimal logical case benefiting from the advanced knowledge of half of the information about the solution.
-Also note that, however, this is just the room for the quantum speedup.
-The quantum algorithm can fully benefit of it, or partly, or not at all.
-Now we go to the evolution of the universe under the fundamental physical laws and the fundamental constants.
-The version of the Anthropic Principle we are dealing with is about the fact that the least change of the value (for values) that the fundamental physical constants have in the universe we are in would have produced a relatively trivial universe unable to develop life.
-This can be formulated as follows.
-We should think of a plurality of universes, all with the same fundamental laws and each with its own value of the fundamental constants.
-We consider any one of these universes, in the mutually exclusive or of all the possible universes.
-The Fine-tuned Universe version of the Anthropic Principle says that its evolution under the same fundamental physical laws and its own value of the physical constants, builds up a (say) one-to-one correlation between the value of the fundamental constants since the beginning of time and the “value of life” (in particular, “existence of sentient life” and “absence of life”).
-Under the quantum cosmological assumption, the evolution of the universe toward the value of life becomes unitary quantum and the one-to-one correlation becomes quantum.
-Under the quantum cosmological reading, the plurality of universes becomes their quantum superposition and the selection of any one universe in the mutually exclusive or of all universes can be performed by an initial measurement (see what follows) of the value of the fundamental constants in the quantum superposition of all the possible universes.
-This is formally identical to a quantum algorithm (see the above extended quantum description of it).
-Therefore, as we shall see, it can occur with quantum speedup, in the sense that there is certainly the room for its occurring with quantum speedup.
-Correspondingly, it can be teleological in character.
-By the way, the notion of measurement requires that of an observer, and, no matter the value of the fundamental constants, at the beginning of time there could have been none.
-However, we can do without this initial measurement.
-There would then be a unitary evolution of the quantum superposition of all possible universes.
-In the universe we are in, and at the time of today, there would be the sentient observer who measures the value of life also selecting the corresponding value of the fundamental constants.
-According to the version of the Anthropic Principle we are dealing with, at the time of today these two observables would already be maximally entangled.
-Since the value of the fundamental constants never changes along the unitary evolution of the universe, all is as if it were selected at the beginning of time.
-Summarizing, in the universe we are in, its evolution toward life can have occurred with quantum speedup (in the sense that there is certainly the room for this) and thus have been teleological in character.
-We shall also see that the same would consequently hold for the evolution of life on Earth (in principle, the trace over all the rest of that of the universe toward life) and the evolutions of species.
-By the way, all the above, the quantum cosmological assumption and the consequent reading of the Fine-tuned Universe version of the Anthropic Principle, can, of course, be considered arbitrary.
-We would argue that it is anyway a plausible possibility, thus very different from the current notion that a teleological evolution of life is an absurdity.
-Of course, if taken seriously, the above would also be highly destabilizing.
-Besides providing a unified quantitative explanation of the quantum speedups, the present result would change the very foundations of the natural sciences, restoring what could be an enlightening relation between evolution and functional structure of the living.
-Again, our purpose here is showing that quantum algorithms are certainly teleological evolution, and their distance from the evolutions of life could even be zeroed under a plausible way of viewing macroscopic reality, i.e.
-the quantum cosmological assumption.
-Also note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science.
-In the present theory of the quantum computational speedup, they could acquire a physical basis.
+However, interestingly, the very context we are dealing with suggests a way out.
+The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11].
+Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such, it can occur with quantum speedup and thus be teleological in character.
+Let us describe how things should go.
+The Fine-tuned universe version of the Anthropic Principle is about the very relevant fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
+Under the quantum cosmological assumption, this version of the Principle can be formalized as follows.
+We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
+Now, always under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental physical laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental physical constants, and each with its own value of the latter.
+Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”.
+The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life.
+Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life; in the other universes, where this value is different, there is absence of life (theory says that, the absence of life in the other universes today is already definitive).
+For what we have said about the generalization of the quantum speedup from quantum algorithms to the building up of quantum correlation, the evolution of the universe toward the value of life can occur with quantum speedup and, therefore, be teleological in character.
+The “can occur” is in the sense that there is room for its occurring with quantum speedup.
+Of course, in the universe we are in, the possible quantum speedup and the consequent teleological character is for the evolution of the universe toward life.
+We will see that this possibility would be inherited by the evolution of life on Earth, in principle, the trace over all the rest of the evolution of the universe toward life.
+Then, for consistency, it would be inherited by the evolutions of species.
+Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered as plausible as the quantum cosmological assumption.
+However, note that all this is quite different from the present widespread belief that the notion of teleological evolution of the living is an absurdity.
+Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science.
+In the present theory of the quantum computational speedup, they would acquire a physical basis.
+Of course, this would not touch on the immense practical value of the mechanistic and reductionist view.
+That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe obviously holds only in principle and does not give what the classical description gives.
+However, this might pinpoint a direction of the research.
+At the theoretical level, it could be interesting to study how this “tracing over all the rest” conciliates a holistic and teleological view in the longer term with a mechanistic and reductionist view in the shorter term.
+This might even be a further perspective under which to explore the unification between classical and quantum mechanics.
 2 Time-symmetric explanation of the quantum speedup
-The present way of thinking about quantum algorithms and their quantum speedups has radical differences from the mainstream one.
-In this section, we provide it at a conceptual level, but always making a point-by-point reference to the formalizations, all postponed in the Appendix.
+We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality.
+Since this is a way of thinking quantum computation radically different from the usual one, we provide it first at a conceptual level.
+All formalizations are postponed in the Appendix.
 The section is divided into subsections corresponding to the various arguments.
-Let us first acknowledge the fact that the discovery of the quantum speedup has been one of the fundamental discoveries of the past century, which, together with quantum communication, has created an entirely new branch of physics.
-However, having said this, for the sake of clarity, we will be completely frank in criticizing what we consider to be the shortcomings of the current way of viewing quantum computation.
-2.1 On the explanation of the quantum speedup
-Let us begin with the advanced knowledge rule, which we repeat here for convenience.
+Let us first acknowledge the fact that that of the quantum computational speedup has been one of the fundamental discoveries of the past century.
+Together with that of quantum communication, it has created an entirely new branch of physics.
+This said, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
+2.1 On the explanation of the quantum computational speedup
+We repeat here for convenience the advanced knowledge rule.
 In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
 The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
-Of course, this rule offers a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information.
+Of course, this rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information.
 Let us discuss this point.
-To the best of our knowledge, the absence of such an explanation has never been acknowledged by the scientific community.
+To the best of our knowledge, the absence of a unified quantitative explanation has never been acknowledged by the scientific community.
 People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition.
@@
 This is naturally our opinion.
-However, as a matter of fact, outside our (destabilizing) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
+However, as a matter of fact, outside our (unconventional) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
+Of course, as already noted, if the only way of providing the unified quantitative explanation in question is by resorting to retrocausality and finalism, the de facto rejection of these notions by the scientific community might have played a role.
 Neither can be considered an explanation the fact that each quantum algorithm naturally explains its quantum speedup.
@@
 Obviously, unification is a fundamental objective of any description of physical phenomena.
-In our view, for what concerns the current mainstream explanation of the quantum speedup, this is the case of crying out: the emperor is naked!
-Our emphasis on this is because it could balance the high unconventionality of our proposal.
+For what concerns the current mainstream explanation of the quantum speedup, let us say, it is the case of crying out that the emperor is naked.
+This emphasis is to underline the absence of any recognition of the fact that the current view of quantum computation has a shortcoming.
 Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation.
-However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against a deeply rooted de facto ban on retrocausality.
-In this work, we should have demonstrated it, indeed in a surprisingly simple way.
-As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the destabilizing character of the rule.
+However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against the deeply rooted de facto ban on retrocausality.
+In this work, we should have demonstrated it, moreover, in a surprisingly simple way.
+As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the unconventional character of the rule.
 2.2 First demonstration of the advanced knoledge rule
-In any case, in order to demonstrate the advanced knowledge rule, we should first amend the current quantum description of quantum algorithms: as it is, it would completely obscure both demonstrations.
-The quantum description in question is limited to the unitary computation of the solution, where the quantum speedup resides, and the measurement of the latter.
-Of course, a canonical quantum description of a quantum process must consist of initial measurement, unitary evolution, and final measurement.
-The current quantum description is obviously focused on the central problem of how to achieve a quantum speedup with respect to to what is logically possible.
-By the way, the explanation of the quantum speedup we will be dealing with completely ignores the way in which it is reached, it is limited to the fact that it is reached.
-Anyway, the above shortcoming is simply remedied: we should add to the current quantum description of the process of solving the problem that of the upstream process of setting the problem.
-An initial measurement in the quantum superposition (or, indifferently, mixture) of all the possible problem-settings, selects one of them at random [^2] The problem-setter could unitarily change it at will; this, here, would change nothing., the unitary computation and the final measurement of the corresponding solution follow—for the formalization of this, see Section (1) of the Appendix.
-We would recommend to look at it; it is trivial but introduces all the definitions that will be used in the subsequent formalizations.
-The quantum description of the quantum algorithm defined until now is with respect to the customary observer external to the quantum process.
-To present ends, we should relativize it with respect to the problem-solver, who computes the solution of the problem.
+We have already given in the Inroduction the quantum description of quantum algorithms extended to the process of solving the problem.
+An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^1] The problem-setter could unitarily change it at will; this, here, would change nothing..
+The unitary computation and the final measurement of the corresponding solution follow.
+For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
+Now, it should also be noted that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process.
+To present ends, we should also relativize it with respect to the problem-solver, who computes the solution of the problem.
 Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the quantum description of the quantum algorithm with respect to her.
-From her, the problem-setting selected by the initial measurement should be concealed.
-As we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action.
+Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action.
+Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easily obtained.
+It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of the unitary evolution.
 This concealment is obtained by postponing to the end of her unitary computation of the solution the projection of the quantum state associated with the initial measurement; see Section (2) of the Appendix.
 Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”.
-It is ordinary because it is still the customary “causal” description of mainstream quantum mechanics—where causality always goes in the same time-direction.
-The first demonstration of the advanced knowledge rule simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
+It is ordinary because it is still the customary “causal” quantum description of mainstream quantum mechanics—where causality always goes in the same time-direction.
+We can go now to the first demonstration of the advanced knowledge rule.
+It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
 By the way, involving only the measurements of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
@@
 (iv) It is optimal in character; this, of course, constitutes an important reference.
-We show how a classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule.
-For the moment, we consider the simplest instance of Grover algorithm.
-As we shall soon see, everything immediately generalizes to the full fledged Grover algorithm and most general forms of quantum correlation (in quantum algorithms, between the problem-setting and the solution of the problem).
+By the way, note that the notion of searching for something and finding it is most universal and this universality is completely represented by Grover algorithm.
+We show that the classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule.
+We begin with the simplest instance of Grover algorithm.
+As we shall soon see, everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
 The problem to be solved is as follows.
@@
 In other words, here, classical logic should not be confused with classical physics, which might be a temptation in the present case.
-Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer.
+Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way, one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer.
 Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution?
@@
 Note that we are dealing with a causal loop.
-Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings).
+Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings); the fewest is the number needed in the optimal logical case benefiting of half of the information about the solution.
 There is a causal loop for each way of taking half of the information about the solution.
@@
 As is well known, the number of drawer openings needed by the (optimal) Grover algorithm is on the order of $\sqrt{N}$ against the order of $N$ of the best classical case, hence its quadratic quantum speedup.
-This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting of this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers.
+This is logically equivalent to knowing in advance, in a mutually exclusive way, that the ball is in one of the possible tuples of $\sqrt{N}$ drawers and benefiting from this knowledge to locate the ball in the best logical way by opening the order of $\sqrt{N}$ drawers.
 Of course, this is what the optimal Grover algorithm does.
@@
 Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers.
-In still equivalent terms, the dimension of the Hilbert in which the search for the solution occurs reduces from $N$ to $\sqrt{N}$.
+In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $N$ to $\sqrt{N}$.
 Holding for Grover algorithm, the above retrocausal explanation of the quantum speedup must hold for the process of solving with a quadratic quantum speedup any NP problem.
@@
 Of course, for the present ends, we are just interested in having a quantum speedup with respect to what is classically possible.
-We shall see further generalizations in the following of this section.
+We shall see that there are further generalizations in the following of this section.
 In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.
@@
 For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm.
-We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of a quantum algorithm, all the information is selected by the initial measurement.
+We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of the quantum algorithm, all the information is selected by the initial measurement.
 Naturally, there is a plurality of possible ways of evenly sharing: we should take their quantum superposition.
 In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement.
-That performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement.
-As we can see in Section (3) of the Appendix, in this new initial measurement outcome it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps.
-This is the physical correspondent of one of the logical causal loops of the advanced knowledge rule.
+Time-symmetrically, that performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement.
+This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum algorithm.
+As we can see in Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum algorithm it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps.
+This is the physical correspondent of one of the logical causal loops whose mutually exclusive or is logically implied by the advanced knowledge rule.
 Eventually, we should take the quantum superposition of all the time-symmetrization instances.
@@
 Of course, there would be no reason why a measurement selected more information than the other.
-As we have just seen, this time symmetrization generates the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule.
+As we have just seen, this time symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup—the physical counterpart of the advanced knowledge rule.
 This is indeed the second demonstration of the advanced knowledge rule.
@@
 In principle, classical logic should also go along with the coexistence of causality and retrocausality..
-This said, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path.
+Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path.
 This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
@@
 We should think that the classical logic description of the sequence of generations of quantum state superpositions and interferences that constitutes the unitary computation of the solution, in the end, simplifies into a description that implies the advanced knowledge rule.
-Summarizing, at the root of the retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it.
+Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it.
 By the way, note that this is not an idle speculation, as might even seem.
-Indeed, we are dealing with a retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
+Indeed, we are dealing with a logical form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
 2.6 Generalization
@@
 However, the possibility of occurring with quadratic quantum speedup holds for any unitary evolution that responds to the rule in question and builds up one-to-one quantum correlation between two initially unrelated observables.
-Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says.
-Of course, in the case that we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
+Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says—see the Appendix.
+Of course, in the case where we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
 The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality.
@@
 With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
-The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops which locally explain it and play the role of the hidden variables envisaged by Einstein et al [9].
+The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15].
 Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement—see [5].
@@
 In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it.
-Of course, it also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
-This ends the first part of the work.
-The second is about the consequences of the advanced knowledge rule or the quantum correlation time-symmetrization rule.
-These consequences are both widespread and destabilizing: indeed, the time-symmetric form of retrocausality implicit in these rules yields a different way of viewing physical reality.
-The consequence just seen, demonstrating the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [9], is destabilizing for the following reason.
-After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [10] and Ekert’s quantum cryptography based on it [11], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
-Another destabilizing consequence of the advanced knowledge rule is that, according to it, quantum algorithms are teleological evolutions, whereas, currently, science claims that teleological evolutions are without physical basis.
-An even more destabilizing consequence is that the physical basis of the teleological character of quantum algorithms, under the quantum cosmological assumption, applies as well to the evolutions of the living.
-Of course, the teleological notion was once applied by natural philosophy to the evolution of life; by natural philosophy, it is also applied today [12].
-As already mentioned, with the advent of modern science, it was dismissed for the alleged absence of a physical basis.
-For its finalistic character, it has also been ascribed to the category of beliefs.
-The next subsection is dedicated to the teleological notion under the present theory of the quantum computational speedup.
-2.7 The physical basis of teleological evolutions
-Let us repeat another time the advanced knowledge rule, which naturally implies the teleological character of quantum algorithms.
-In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
-The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
-We have demonstrated this rule in two redundant ways.
-The first, by describing the quantum speedup in a classical logic way, the second by a time-symmetrization of the ordinary quantum description of the quantum algorithm which is mandatory to satisfy the atemporal symmetry implicit in it: we have seen that this time-symmetrization generates the physical correspondent of the advanced knowledge rule.
-With this, of course, we have also demonstrated the teleological character of quantum algorithms: indeed, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
-We have also seen that, under the quantum cosmological assumption that the universe evolves in a unitary quantum way, and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the value of life acquires the form of a quantum search algorithm.
-As such, it can occur with a quadratic quantum speedup.
-The “can occur” is in the sense that there is certainly room for a quantum speedup, but the unitary evolution must benefit of it; it should occur with a just suboptimal quantum speedup if the search for the value of life is random [21] .
-We go now to the fact that this possible teleological character would be inherited by the evolution of life on Earth and the evolution of species.
-Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life applies to the evolution of the universe toward life.
+Of course, this also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
+By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [15] might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
+2.7 About the teleological character of the evolutions of the living
+We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolution of the universe toward the value of life.
+We repeat here for convenience the reason for its possible quantum speedup and consequent teleological character.
+Under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental constants, and each with its own value of the latter.
+Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”.
+The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life.
+Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life, in the other universes where this value is different there is absence of life.
+We have seen in the Generalization subsection that any such building up of quantum correlation can always occur with quantum speedup in the sense that there is certainly the room for it.
+It can consequently be teleological in character.
+Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life holds for the evolution of the universe toward life.
 Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth.
-In principle, the latter evolution is derivable from the former evolution by tracing out all the rest.
+In principle, the latter evolution can be derived from the former evolution by tracing all the rest.
 Therefore, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life.
@@
 In fact, the latter resources comprise the former.
-In summary, the evolution of life on Earth could occur with quantum speedup (in the sense that there is certainly the room for that) and consequently be teleological in character; it should actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random.
+In summary, under the quantum cosmological assumption, the evolution of life on Earth can occur with quantum speedup and consequently be teleological in character; it might actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random.
 Then, just for consistency, the evolutions of species would also be the same.
@@
 See our preprint [6] for a more comprehensive discussion of the teleological notion.
-All this is obviously destabilizing, it is a way of viewing the macroscopic reality that goes against the modern way of viewing it, and revamps ancient natural philosophy views which were ascribed by science to completely unscientific beliefs.
-Let us then discuss more extensively this contrast between the scientific and the natural philosophy view of the teleological notion.
-The notion of teleology as final cause was introduced almost two and a half millennia ago by Plato and Aristotle to explain the “formation” of the functional structure of the living (Aristotle: teeth were created for the purpose of chewing).
-Of course, at the time, there was not yet the notion of evolution of the living.
-In more recent times, the teleological notion was applied by natural philosophy to the evolutions of the living—by natural philosophy it is also applied today [12].
-It was then dismissed by science for the alleged absence of a physical basis—precisely because it implies a retrocausality deemed to be physically impossible.
-Of course, at the time of classical physics, this was completely justified.
-For its finalistic character, the teleological notion was also ascribed to the category of beliefs.
-It should however be noted that this dismissal was not without regret; indeed, it meant doing without a possibly enlightening relation between the notion of teleological evolution of life and that of the functional structure of living (teeth were created for the purpose of chewing).
-We cite two authoritative objections to this dismissal.
-One is a famous sentence by the evolutionary biologist Haldane [26]:
-Teleology is like mistress to a biologist, he cannot live without her, but is unwilling to be seen with her in public.
-The other is the position of the natural philosopher Nagel [12], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view.
-In our possibly biased opinion, this would be a case in which common sense prevails over a de facto no-retrocausality precept that, according to the present work, would be a prejudice.
+All this is obviously destabilizing; it is a way of viewing the macroscopic reality that goes against the current (scientific) way of viewing it, and revamps the natural philosophy views that were ascribed by science to the unscientific beliefs.
+Of course, the fact that what occurs to us today is affected by the drying up of an area of Africa three millions of years ago is obvious, but that it is also affected by what we will be in three million years’ time, to the current way of thinking is more difficult to digest.
 This also raises the question of how to describe a macroscopic evolution endowed with quantum speedup.
@@
 It is explained by a quantum model, whereas it would be difficult to explain it in a classical way.
-In our case, the canonical way of describing the evolution of life on Earth would be to obtain it from the description of the evolution of the universe toward life by tracing out all the rest.
+In our case, the canonical way of describing the evolution of life on Earth would be to obtain it from the description of the evolution of the universe toward life by tracing-out all the rest.
 Since this obviously holds only in principle, we should also model the macroscopic teleological evolutions of species endowed with quantum speedup in a quantum way.
@@
 This situation might point to a particular object of research.
-It could be interesting to study how the operation of “tracing-over the rest” (at the toy model level) reconciles a teleological behavior in the long term with a classical behavior in the short term.
+It could be interesting to study how the operation of “tracing-out all the rest” (at the toy model level) reconciles the teleological character of an evolution in the longer term with its classical character in the shorter term.
+This could also be a further perspective under which to address the problem of unifying classical and quantum theory.
 Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup.
-According to the present theory, it would be an even more fundamental discovery than currently thought.
-The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [15] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—an attempt to unify relativity and quantum theory.
+According to the present theory, this discovery would be even more fundamental than currently thought.
+The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [19] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of an attempt to unify relativity and quantum theory.
 Then, there was the introduction of the notion of logically reversible computation by Bennett [20] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions.
@@
 Interestingly, at the time, Costa de Beauregard was working for de Broglie, who forbid him to publish that work involving retrocausality for years.
-Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory.
+Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory, which also implied a form of retrocausality.
 See [25] for an account of this episode.
@@
 However, first, none of them concerns the quantum computational speedup and, second, they all apply to very specific physical situations to which they might be confined.
-On the contrary, the present time-symmetric form of retrocausality should be sort of universal for quantum correlations.
+On the contrary, the present time-symmetric form of retrocausality is sort of universal for quantum correlations.
 The possible relation of those works to the present work, as things are now, would not be evident.
-Of course, it could be the object of further investigation.
+Of course, it could be the object of study.
 4 Conclusions
@@
 More specifically, since causality could well go backward in time along a unitary quantum evolution that is both objectively and ideally time-reversible, we should say that, in the quantum world, the scientific tradition amounts to asserting the mutual exclusivity between the two possible time directions of causality.
-Let us first address the reliability of the work, a key issue because of its unconventionality.
+Let us first address the reliability of the work, a particular issue because of its unconventionality.
 That the classical logic description of the quantum speedup must resort to retrocausality should be obvious.
@@
 The other is that all the results obtained under the assumption that quantum algorithms can be described by classical logic are also obtained by time-symmetrizing their ordinary quantum description; this, in turn, is mandatory to satisfy the atemporal symmetry between the simultaneous measurements of two maximally entangled observables implicit in the ordinary quantum description of quantum algorithms.
-We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences (or right for them).
+We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences (or, according to tastes, even right for them).
 The other side of the coin of the destabilizing character of the work is, of course, its revolutionary character, always due to the retrocausality implicit in it.
@@
 (ii) It theoretically confirms the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it.
-(iii) It highlights the teleological character of quantum algorithms; by the way, given the negative view that the teleological notion has in science today, and the scope of this notion, this should be the most revolutionary result of all.
-About the plausible things: It shows that, under a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could apply as well to the evolutions of the living.
-Let us eventually note that all this should change our way of seeing physical reality.
-Of course, for this, it should be sufficient the revamping of the teleological notion.
-Why now?
-Naturally, after the relatively recent discovery of the quantum computational speedup.
+(iii) It highlights the teleological character of quantum algorithms; by the way, given the negative view that the teleological notion has today in science and the scope of this notion, this should be the most revolutionary result of all.
+What is just plausible: It shows that, under a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
+Let us eventually note that all this would change our way of viewing physical reality.
+Curiously, it would be a case where the progress of science, the advent of quantum mechanics, brought us back to a way of viewing physical reality that preceded science.
 Appendix
@@
 As anticipated in the Introduction, the ordinary quantum description of Grover algorithm is incomplete and is completed by time-symmetrizing it.
-We should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes evenly shares between the initial and final measurements.
-The half information selected by the initial measurement should propagate forward in time, by $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement.
-Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time, by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$, until replacing the previous outcome of the initial measurement.
+We should require that the selection of the information that specifies the sorted out pair of correlated measurement outcomes, the problem-setting and the corresponding solution, evenly shares between the initial measurement of the problem setting and the final measurement of the solution.
+The half information selected by the initial measurement should propagate forward in time by the unitary computation of the solution $\hat{\mathbf{U}}_{1,2}$ until becoming the state immediately before the final measurement.
+Time-symmetrically, the complementary selection performed by the final measurement should propagate backward in time by the inverse of the computation of the solution $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until replacing the previous outcome of the initial measurement.
 We consider the time-symmetrization instance where the initial measurement of $\hat{B}$ reduces to the measurement of $\hat{B}_{l}$, the left bit of the two-bit number contained in register $B$; say that it randomly selects left bit 0.
@@
 Then (in causal order), this final measurement outcome propagates backward in time by $\hat{\mathbf{U}}_{1,2}^{{\dagger}}$ until becoming the definitive outcome of the initial measurement.
-This backward in time propagation is an instance of the time-symmetrized description of the quantum algorithm with respect to the problem-solver.
-As we can see from the diagram, it tells her, before she begins her problem-solving action, one of the possible halves of the information about the solution, namely that the ball is in drawer either 01 or 11: look at the state of register $B$ in the bottom left corner of the table.
-In summary, the causal loop is the zigzag diagram of Table III.
+This backward in time propagation is an instance of the time-symmetrized quantum algorithm.
+As we can see from the diagram, the new outcome of the initial measurement tells the problem-solver, before she begins her problem-solving action, one of the possible halves of the information about the solution, namely that the ball is in drawer either 01 or 11: look at the state of register $B$ in the bottom left corner of the table.
+Note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, this bottom line of Table III diagram (an instance of the time-symmetrized quantum algorithm) mathematically can also be read in the usual left to right way:
+$\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
+We can see that, in each instance of the time-symmetrized quantum algorithm, the computational complexity of the problem to be solved by Alice quadratically reduces from the problem of locating the ball in one of $4$ drawers to that of locating it in $2=\sqrt{4}$ drawers.
+In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces from $4$ basis vectors to $2=\sqrt{4}$ basis vectors.
+In summary, the causal loop generated by the time-symmetrization process is the zigzag diagram of Table III.
 We can see that it is one of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup.
@@
 Through the zig-zag, this state changes into the state of register $B$ in the bottom-left part of the diagram, which tells us that Alice, before beginning her problem-solving action, knows in advance that the ball is in drawer either $01$ or $11$; we are dealing with an effective causal loop that changes its own past.
-She can then solve the problem in an optimal logical way by opening just one of the two drawers, indeed as in the classical logic description of what the quantum algorithms does: look at the bottom line of the diagram in Table III.
-Note that, for the mathematical equivalence between $\Leftarrow\hat{\mathbf{U}}_{1,2}^{{\dagger}}\Leftarrow$ and $\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow$, this bottom line of Table III diagram (mathematically) can also be read in the usual left to right way:
-$\begin{array}[]{ccc}\text{time }t_{1}&t_{1}\rightarrow t_{2}&\text{time }t_{2}\\ &&\\ \begin{array}[]{c}(|01\rangle_{B}+|11\rangle_{B})\\ \left(|00\rangle_{A}+|01\rangle_{A}+|10\rangle_{A}+|11\rangle_{A}\right)\end{array}&\Rightarrow\hat{\mathbf{U}}_{1,2}\Rightarrow&|01\rangle_{B}|01\rangle_{A}+|11\rangle_{B}|11\rangle_{A}\end{array}$ (Table IV)
-We can see that, in each time-symmetrization instance (causal loop) the computational complexity of the problem to be solved by Alice quadratically reduces from the problem of locating the ball in one of $4$ drawers to that of locating it in $2=\sqrt{4}$ drawers.
-In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place quadratically reduces from $4$ basis vectors to $2=\sqrt{4}$ basis vectors.
+She can then solve the problem in an optimal logical way by opening just one of the two drawers, indeed as in the classical logic description of what the quantum algorithms does.
+Also note that table III diagram goes along with the fact that the unitary part of the quantum algorithm, $\hat{\mathbf{U}}_{1,2}$, takes the opening just one drawer, but does not imply it.
+We know this fact from outside Table III diagram: we know that this is what the unitary part of Grover algorithm takes.
+The diagram says that the problem-solver knows in advance that the ball is in a pair of drawers rather than four—bottom left corner of the diagram—and this could allow her to solve the problem by opening a single drawer—bottom line of the diagram of table III or the diagram of table IV.
+The advanced knowledge rule tells us the room for the quantum speedup but does not imply it.
 Eventually, we should take the quantum superposition of all the time-symmetrization instances.

### v3

@@
 1 Introduction
-Let us briefly examine the very long history of the teleological notion—attraction by a final cause.
-Plato and Aristotle used it to explain the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution.
+Let us briefly retrace the two and a half millennia history of the teleological notion: attraction exercised by a final cause.
+Plato and Aristotle applied it the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution.
 With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy.
-It was eventually dismissed by science for the alleged absence of a physical basis.
-All this was before the advent of quantum mechanics.
+It was eventually dismissed by science (it was classical science) for the alleged absence of a physical basis, even ascribed, for its finalistic character, to the category of beliefs.
+All this was before the advent of quantum mechanics, but has remained almost unchanged with it.
 Today, the situation of the teleological notion is somewhat complex.
-To the best of our knowledge, outside the present work, there are no physical researches on it.
+However, to the best of our knowledge, outside the present work, in physics, there are no researches on it.
 Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
 In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science.
-This is because, in the first place, quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
-Since showing this will be surprisingly simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
+This is because quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
+Since showing this will be simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
 We, with coauthors, already went against this position in the evolutionary, retrocausal, approach to the quantum computational speedup [1, 2, 3, 4, 5, 6].
@@
 We can therefore say that, according to this rule, the quantum speedup is explained by a mutually exclusive or of causal loops, one for each possible way of taking half of the information about the solution.
-By the way, here and in the following by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible.
+By the way, here and in the following, by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible.
 Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
@@
 The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem).
+We will see further on that the physical definition of teleological evolution must be that of an evolution between an initial and a final measurement outcome.
 Until now, this advanced knowledge rule was, so to speak, an empirical rule shown to be satisfied by all major quantum algorithms.
@@
 Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
-As a matter of fact, we have provided two redundant demonstrations of this rule, what should not be despised given its destabilizing character. Let us outline them.
+As a matter of fact, we will provide two redundant demonstrations of this rule, which should not be despised given its destabilizing character. Let us outline them.
 One is an application of classical logic to what the quantum algorithms does: providing a quantum speedup with respect to what is classically possible.
-By the way, we will have to use a quantum description of the quantum algorithms that is not their customary quantum description.
+By the way, we will have to use a quantum description of the quantum algorithms that is not the customary one.
 The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution.
 It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement.
-We complete the customary quantum description by adding to the description of the process of solving the problem that of the upstream process of setting the problem.
+We complete this customary quantum description by adding, to the description of the process of solving the problem, that of the upstream process of setting the problem.
 The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow.
 Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic.
-As we shall see in the next section, the classical logic description of the quantum speedup is indeed the advanced knowledge rule.
+Having said this, the outline of the first demonstration can be very short.
+As we shall see in the next section, Under the present more complete quantum description of the quantum algorithm, the classical logic description of the quantum speedup turns out to be indeed the advanced knowledge rule.
 Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete.
 Indeed, it is simply completed by time-symmetrizing it.
-We shall see that the time-symmetrization of the quantum description yields the quantum superposition of the causal loops involved in the advanced knowledge rule.
-In other words, the time-symmetrization of the quantum description of the quantum algorithm generates the physical correspondent of the advanced knowledge rule.
-The second demonstration of the advanced knowledge rule simply comes from the fact that this time-symmetrization of the quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far.
-As we shall see, it is to satisfy an atemporal symmetry implicit in the quantum description.
-Of course, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—it is a demonstration of the rule.
-Of course, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
-A further step, important for what concerns the possible teleological character of the evolutions of the living, is that all the above generalizes to quantum correlation.
-For simplicity, we limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty.
-We can then say that the quantum algorithm builds up one-to-one correlation between two initially unrelated observables (the problem-setting and the solution of the problem) without ever changing the eigenvalue of the observable selected by the initial measurement (the problem-setting).
-We shall show that any building up of quantum correlation of this kind (not only that in quantum algorithms) can occur with quantum speedup and therefore be teleological in character.
-Note that the “can be” means that there is certainly room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation.
+We shall see that the time-symmetrization of the more complete quantum description of the quantum algorithm generates the quantum superposition of the causal loops involved in the advanced knowledge rule.
+The second demonstration of the advanced knowledge rule simply comes from the fact that the time-symmetrization of the more complete quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far.
+As we shall see, it is to satisfy an atemporal symmetry implicit in this description.
+As we have seen, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—we have thus a demonstration of the rule.
+Naturally, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
+A further step, important for the possibility of extending the teleological character of quantum algorithms to the evolutions of the living, is the following generalization.
+For simplicity, let us limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty.
+We can then say that the quantum algorithm unitarily builds up one-to-one correlation between two initially unrelated observables—the problem-setting and the solution of the problem—without ever changing the eigenvalue of the observable selected by the initial measurement—the problem-setting.
+We shall show that any building up of quantum correlation of this kind, also outside quantum computation, can occur with quantum speedup and therefore be teleological in character.
+Note that the “can occur” means that there is certainly the room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation.
 By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
-With this, we can go to the evolutions of the living for which the teleological notion was originally conceived.
+We go now to the evolutions of the living for which the teleological notion was conceived.
 First, let us note that the dismissal of the teleological notion was not without (scientific) regret; indeed, it meant doing without a possibly enlightening relation between the notion of evolution of life and that of the functional structure of the living.
@@
 The latter have a duration of billions of years and can have the size of the Earth.
-However, interestingly, the very context we are dealing with suggests a way out.
+However, the very context we are dealing with suggests a way out.
 The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11].
-Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) acquires the form of a quantum algorithm and, as such, it can occur with quantum speedup and thus be teleological in character.
+Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) could acquire the form of a quantum algorithm and, as such, it could occur with quantum speedup and thus be teleological in character.
 Let us describe how things should go.
-The Fine-tuned universe version of the Anthropic Principle is about the very relevant fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
+The Fine-tuned universe version of the Anthropic Principle is about the extraordinary theoretical fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
 Under the quantum cosmological assumption, this version of the Principle can be formalized as follows.
 We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
-Now, always under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental physical laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental physical constants, and each with its own value of the latter.
-Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”.
-The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life.
-Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life; in the other universes, where this value is different, there is absence of life (theory says that, the absence of life in the other universes today is already definitive).
-For what we have said about the generalization of the quantum speedup from quantum algorithms to the building up of quantum correlation, the evolution of the universe toward the value of life can occur with quantum speedup and, therefore, be teleological in character.
-The “can occur” is in the sense that there is room for its occurring with quantum speedup.
-Of course, in the universe we are in, the possible quantum speedup and the consequent teleological character is for the evolution of the universe toward life.
-We will see that this possibility would be inherited by the evolution of life on Earth, in principle, the trace over all the rest of the evolution of the universe toward life.
-Then, for consistency, it would be inherited by the evolutions of species.
-Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered as plausible as the quantum cosmological assumption.
-However, note that all this is quite different from the present widespread belief that the notion of teleological evolution of the living is an absurdity.
+Of course, at the beginning of time, these two observables must be unrelated: no matter the value of the fundamental constant, the value of life is “absence of life”.
+At the time of today, the unitary evolution of the universe toward the value of life has built up one-to-one quantum correlation between the value of the fundamental constants since the beginning of time and the value of life, of course, without ever changing the former value.
+We can therefore say that the evolution of the universe toward the value of life can occur with quantum speedup and thus be teleological in character; the “can occur” is in the sense that there is certainly the room for its occurring with quantum speedup.
+Of course, whether the quantum speedup is achieved depends on the character of the unitary evolution.
+For what was said about quantum random search algorithms, we might think that, if the search for the value of life by the unitary evolution of the universe is random, it should occur with a quantum speedup.
+In the universe we are in, the evolution of the universe toward the value of life becomes evolution of the universe toward life; of course, all the above also applies to the latter evolution.
+We shall see that, as a consequence, the evolution of life on Earth, in principle the trace over all the rest of the evolution of the universe toward life, would inherit the possible teleological character of the evolution of the universe toward life.
+For consistency, the same would apply to the evolution of species.
+Of course, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption.
+Anyway, all this should be quite different from the present widespread belief that the notion of a teleological evolution of the living is an absurdity.
+At least, it should cast a doubt on the present position of science.
 Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science.
@@
 Of course, this would not touch on the immense practical value of the mechanistic and reductionist view.
-That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe obviously holds only in principle and does not give what the classical description gives.
-However, this might pinpoint a direction of the research.
-At the theoretical level, it could be interesting to study how this “tracing over all the rest” conciliates a holistic and teleological view in the longer term with a mechanistic and reductionist view in the shorter term.
-This might even be a further perspective under which to explore the unification between classical and quantum mechanics.
+That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe, obviously, holds only in principle: in practice, we cannot have such descriptions.
+However, this contrast between the quantum and the classical view might pinpoint a direction of the research.
+At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system in the longer term with a classical character in the shorter term.
+This might even be a further perspective under which to study the unification between quantum and classical mechanics.
 2 Time-symmetric explanation of the quantum speedup
 We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality.
-Since this is a way of thinking quantum computation radically different from the usual one, we provide it first at a conceptual level.
+Since this is a way of thinking quantum computation that is radically different from the usual one, we provide it first at a conceptual level.
 All formalizations are postponed in the Appendix.
@@
 Together with that of quantum communication, it has created an entirely new branch of physics.
-This said, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
+However, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
 2.1 On the explanation of the quantum computational speedup
@@
 Obviously, unification is a fundamental objective of any description of physical phenomena.
-For what concerns the current mainstream explanation of the quantum speedup, let us say, it is the case of crying out that the emperor is naked.
-This emphasis is to underline the absence of any recognition of the fact that the current view of quantum computation has a shortcoming.
+Summing up, in our judgment, for what concerns the current mainstream explanation of the quantum speedup, it is the case of saying that the emperor is naked.
 Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation.
-However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it was not demonstrated; second, it goes against the deeply rooted de facto ban on retrocausality.
+However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it has not been demonstrated; second, it goes against the de facto ban on retrocausality.
 In this work, we should have demonstrated it, moreover, in a surprisingly simple way.
-As a matter of fact, we shall be dealing with two independent and redundant demonstrations, which is not to be despised because of the unconventional character of the rule.
+As a matter of fact, we shall be dealing with two independent demonstrations, which is not to be despised because of the destabilizing character of the rule.
 2.2 First demonstration of the advanced knoledge rule
-We have already given in the Inroduction the quantum description of quantum algorithms extended to the process of solving the problem.
+We have already provided in the Introduction the quantum description of quantum algorithms extended to the process of solving the problem.
 An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^1] The problem-setter could unitarily change it at will; this, here, would change nothing..
@@
 For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
-Now, it should also be noted that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process.
-To present ends, we should also relativize it with respect to the problem-solver, who computes the solution of the problem.
-Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the quantum description of the quantum algorithm with respect to her.
-Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would also tell her the solution of the problem before she begins her problem-solving action.
-Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easily obtained.
-It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of the unitary evolution.
-This concealment is obtained by postponing to the end of her unitary computation of the solution the projection of the quantum state associated with the initial measurement; see Section (2) of the Appendix.
+Now, we should note that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process.
+To present ends, we should relativize it with respect to the problem-solver who, for the fact of computing the solution of the problem, is part of the process.
+Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the completed quantum description of the quantum algorithm with respect to her.
+Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would tell her the solution of the problem before she begins her problem-solving action.
+Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easy.
+It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of her unitary computation of the solution; see Section (2) of the Appendix.
 Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”.
@@
 (iv) It is optimal in character; this, of course, constitutes an important reference.
-By the way, note that the notion of searching for something and finding it is most universal and this universality is completely represented by Grover algorithm.
-We show that the classical logic description of the quantum speedup delivered by Grover algorithm implies the advanced knowledge rule.
+By the way, note that the notion of searching for something and finding it is most universal and that this universality is completely represented by Grover algorithm.
+We show that the classical logic description of the quantum speedup delivered by Grover algorithm very directly implies the advanced knowledge rule.
 We begin with the simplest instance of Grover algorithm.
@@
 Note that we are dealing with a causal loop.
-Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution with fewer computation steps (drawer openings); the fewest is the number needed in the optimal logical case benefiting of half of the information about the solution.
+Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future (that the ball is in two rather than four drawers) and can use this knowledge to produce the solution by opening just one drawer.
 There is a causal loop for each way of taking half of the information about the solution.
@@
 Also in this case, using the advanced knowledge of half of the information about the solution in an optimal logical way yields the optimal quantum speedup of Grover algorithm.
+This is, of course, the advanced knowledge rule for the full fledged Grover algorithm.
 Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers.
@@
 The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete.
-Of course, to bring it in line with the classical logic description, it should be completed.
+Of course, to bring it in line with their classical logic description, it should be completed.
 It turns out that this is simply done by time-symmetrizing it.
@@
 Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^2] By the way, classical logic is neutral for what concerns the notion of causality.
-Customarily associating it with a causality that always goes in the same time-direction must be consequent to the de facto no-retrocausality precept.
-In principle, classical logic should also go along with the coexistence of causality and retrocausality..
+Associating it with a causality that always goes in the same time-direction, as customary, must be consequent to the de facto no-retrocausality precept.
+Classical logic should also go along with the coexistence of causality and retrocausality..
 Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path.
@@
 We should think that the classical logic description of the sequence of generations of quantum state superpositions and interferences that constitutes the unitary computation of the solution, in the end, simplifies into a description that implies the advanced knowledge rule.
-Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fact that the classical logic description of the generation of a quantum state superposition (one of the pillars of quantum mechanics) necessarily involves it.
-By the way, note that this is not an idle speculation, as might even seem.
-Indeed, we are dealing with a logical form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
+Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fundamental notion of quantum state superposition, the fact that the classical logic description of its generation necessarily involves retrocausality.
+By the way, note that this is not an idle speculation as it might even seem.
+Indeed, we are dealing with a form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
 2.6 Generalization
 In the second demonstration of the advanced knowledge rule, we have shown that a unitary evolution between two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it.
-Note that we never needed to look inside that unitary evolution $\hat{\mathbf{U}}_{1,2}$: see Tables I through IV of the Appendix.
-Therefore, this time-symmetrization does not only hold for quantum algorithms, it lends itself to an important generalization:
+Note that we never needed to look inside the unitary evolution, for example $\hat{\mathbf{U}}_{1,2}$ of Tables I through IV of the Appendix.
+Thus, the need for this time-symmetrization does not only hold for quantum algorithms; there is the following generalization:
 The ordinary quantum description of any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it.
@@
 With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
-The time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15].
-Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement—see [5].
-Note that this is what Einstein et al required from their hidden variables.
-That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process.
+As shown in [5], the time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15].
+Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement.
+Note that this is exactly what Einstein et al required from their hidden variables.
+That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process equivalent in this case.
 In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it.
 Of course, this also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
-By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al in their EPR paper [15] might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
+By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
+By the way, if the only possible hidden variables are retrocausal in character as the ones we are dealing with, this dismissal can be ascribed to the already mentioned de facto rejection of retrocausality and finalism (the form it has in philosophy, different from the present trivial form).
 2.7 About the teleological character of the evolutions of the living
-We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms also applies to the evolution of the universe toward the value of life.
-We repeat here for convenience the reason for its possible quantum speedup and consequent teleological character.
-Under the quantum cosmological assumption, we should consider a quantum superposition of all the possible universes, all with the same fundamental laws (thus, the same unitary transformation from the beginning of time until today) but for the value of the fundamental constants, and each with its own value of the latter.
-Initially, the value of the fundamental constants and that of life are two independent observables: indeed, no matter the value of the fundamental constants, the value of life is “absence of life”.
-The subsequent unitary evolution of this quantum superposition, from the beginning of time until today, builds up a one-to-one correlation between the value of the fundamental constants and the value of life.
-Today, in the universe we are in, the value of the fundamental constants is that which produces sentient life, in the other universes where this value is different there is absence of life.
-We have seen in the Generalization subsection that any such building up of quantum correlation can always occur with quantum speedup in the sense that there is certainly the room for it.
-It can consequently be teleological in character.
-Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value of life holds for the evolution of the universe toward life.
+We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could also apply to the evolution of the universe toward the value of life.
+This section expands on some facts that, for compactness, were just mentioned in the Introduction.
+Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value becomes the possible teleological character of the evolution of the universe toward life.
 Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth.
-In principle, the latter evolution can be derived from the former evolution by tracing all the rest.
-Therefore, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life.
+In principle, the latter evolution can be derived from the evolution of the universe toward life by tracing all the rest.
+Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life.
 Then, first, evolution times must be the same for the evolution of life on Earth and its image in the evolution of the universe.
@@
 In fact, the latter resources comprise the former.
-In summary, under the quantum cosmological assumption, the evolution of life on Earth can occur with quantum speedup and consequently be teleological in character; it might actually occur with quantum speedup if the search for the value of life by the evolution of the universe toward life were random.
+In summary, under the quantum cosmological assumption, the evolution of life on Earth could occur with quantum speedup (in the sense that there would be the room for this) and consequently be teleological in character; it might actually occur with quantum speedup if the search for life by the evolution of the universe toward life were random.
 Then, just for consistency, the evolutions of species would also be the same.
-Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would mean for the notion of the evolutions of species.
-It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs that did not fly.
+Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; it will be very destabilizing with respect to the present way of viewing the evolutions of the living.
+Let us begin with the evolution of species.
+It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs who did not fly.
 Very light, strong but flexible, cave and air-proof, even at that time feathers were ideal for flight, very close to those of the future birds.
@@
 The usual explanation, of course, is causal.
-First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually, feathers turned out to be what was needed to conquer the air.
+First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually feathers turned out to be what was needed to conquer the air.
 All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred.
 See our preprint [6] for a more comprehensive discussion of the teleological notion.
-All this is obviously destabilizing; it is a way of viewing the macroscopic reality that goes against the current (scientific) way of viewing it, and revamps the natural philosophy views that were ascribed by science to the unscientific beliefs.
-Of course, the fact that what occurs to us today is affected by the drying up of an area of Africa three millions of years ago is obvious, but that it is also affected by what we will be in three million years’ time, to the current way of thinking is more difficult to digest.
-This also raises the question of how to describe a macroscopic evolution endowed with quantum speedup.
+All this is obviously destabilizing; it is a way of viewing the time-development of macroscopic reality that goes against the current scientific (causal) way of viewing it.
+Indeed, it revamps the natural philosophy views that were ascribed by science to the category of the unscientific beliefs.
+To the above example of teleological evolution of life, which concerned the evolution of the DNA, we should add another example that concerns the environment of the evolution of life.
+For example, what we humans are today has been affected by the drying up of an area of Africa in which our ancestors lived on trees three million years ago.
+Of course, there is nothing destabilizing in this.
+However, the time-symmetric fact that what we humans are today is also affected by what will concern our descendants in three million years’ time might be very difficult to digest.
+Indeed, it would profoundly change our (causal) way of viewing how things evolve in time.
+Of course, all this is under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle and is thus hypothetical.
+What we would consider certain is the teleological character of quantum algorithms.
+Things could be further explored by investigating, either theoretically or by means of toy models, the character of the “tracing over all the rest”.
+We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing over all the rest.
+We should look for a way of obtaining an evolution of the subsystem that is teleological in the longer term and (about) classical in the shorter term.
+If this were possible, of course, it would go in favor of the above quantum cosmological reading of that version of the Anthropic Principle.
+By the way, it might also be a way of addressing the problem of unifying quantum and classical theory.
+Another question is how to describe a macroscopic evolution endowed with quantum speedup.
 Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
@@
 Indeed, if it were completely autonomous, of course, it would be a unitary quantum evolution itself.
-Incidentally, it might also be interesting to derive the description of the evolution of life on Earth from that of the universe toward life by tracing out all the rest at the level of toy models of the two evolutions.
-Within the present theory, this might highlight significant characters of a possible teleological evolution of life on Earth.
-By the way, of course, the quantum cosmological view of the evolution of the universe would “unify” quantum and classical physics simply by canceling the latter.
-However, naturally, classical physics turns out to work extremely well in describing macroscopic reality.
-Presently, it would have the shortcoming of being unable to describe macroscopic teleological evolutions.
-This situation might point to a particular object of research.
-It could be interesting to study how the operation of “tracing-out all the rest” (at the toy model level) reconciles the teleological character of an evolution in the longer term with its classical character in the shorter term.
-This could also be a further perspective under which to address the problem of unifying classical and quantum theory.
 Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup.
@@
 The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [23] in 1985.
-Note how the notion of the logical reversibility of computation played an essential role in the development of the notion of quantum computation.
-The present time-symmetric quantum description of quantum algorithms could be a further step in this direction.
-Involving a time-symmetric coexistence of the two possible time-directions of causality, it could be said that it exploits the objectively ideal time-reversibility of unitary quantum evolutions in the possibly deepest way.
+Note the importance of the role that the notion of reversibility played in the development of quantum computation.
+The present work could be put on the same path.
+It could be said that the present time-symmetric form of causality exploits the objectively ideal time-reversibility of unitary quantum evolutions in the deepest possible way.
 3 Precedents
 We have seen that applying the quantum correlation time-symmetrization rule to quantum nonlocality provides a local explanation of it. Now, we recall a historical precedent of this.
-It is a work by Costa de Beauregard [24] that also provides a retrocausal explanation of quantum nonlocality.
+It is a work by Costa de Beauregard [24, 25] that also provides a retrocausal explanation of quantum nonlocality.
 In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement.
 Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar.
-Interestingly, at the time, Costa de Beauregard was working for de Broglie, who forbid him to publish that work involving retrocausality for years.
-Ultimately, that ban was withdrawn only after the publication of Wheeler and Feynman’s work on the time-symmetric emitter-absorber theory, which also implied a form of retrocausality.
-See [25] for an account of this episode.
-It is rather significant of the attitude, of even top physicists, toward the retrocausality notion.
 Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [26, 27].
 Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics.
-However, first, none of them concerns the quantum computational speedup and, second, they all apply to very specific physical situations to which they might be confined.
-On the contrary, the present time-symmetric form of retrocausality is sort of universal for quantum correlations.
-The possible relation of those works to the present work, as things are now, would not be evident.
-Of course, it could be the object of study.
+However, none of them concerns the quantum computational speedup.
+Their possible relation to the present work, as things are now, would not be evident.
 4 Conclusions
@@
 That the classical logic description of the quantum speedup must resort to retrocausality should be obvious.
-Indeed, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers.
+Of course, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers.
 To the problem-solver, the only possible source of this information is the solution of the problem she will produce and measure in the future: hence retrocausality.
@@
 The other is that all the results obtained under the assumption that quantum algorithms can be described by classical logic are also obtained by time-symmetrizing their ordinary quantum description; this, in turn, is mandatory to satisfy the atemporal symmetry between the simultaneous measurements of two maximally entangled observables implicit in the ordinary quantum description of quantum algorithms.
-We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences (or, according to tastes, even right for them).
-The other side of the coin of the destabilizing character of the work is, of course, its revolutionary character, always due to the retrocausality implicit in it.
-We should distinguish between what should be certain and what could be just plausible.
-Beginning with what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedup, which, forty years after its discovery, was still missing from mainstream quantum mechanics.
-(ii) It theoretically confirms the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it.
-(iii) It highlights the teleological character of quantum algorithms; by the way, given the negative view that the teleological notion has today in science and the scope of this notion, this should be the most revolutionary result of all.
-What is just plausible: It shows that, under a quantum cosmological interpretation of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
-Let us eventually note that all this would change our way of viewing physical reality.
-Curiously, it would be a case where the progress of science, the advent of quantum mechanics, brought us back to a way of viewing physical reality that preceded science.
+We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences.
+The flip side of the destabilizing character of the work is its innovative character.
+Let us distinguish between what should be certain and what should be plausible.
+About what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics.
+(ii) It provides a theoretical confirmation of the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it.
+(iii) It highlights the teleological character of quantum algorithms; given the negative view that the teleological notion has in today’s science and the scope of this notion, this should be the most important result.
+About what is just plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
 Appendix

### v4

@@
 For consistency, the same would apply to the evolution of species.
-Of course, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption.
+Note that the fact that a teleological evolution is also an evolution endowed with a quantum speedup is a novelty for the teleological notion.
+It could contribute to explaining steps of natural evolutions that are too rapid to be explained by the progressive character of Darwinian evolutions.
+Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption.
 Anyway, all this should be quite different from the present widespread belief that the notion of a teleological evolution of the living is an absurdity.
-At least, it should cast a doubt on the present position of science.
+At least, it should cast a doubt on this position of science.
 Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science.
@@
 Of course, this would not touch on the immense practical value of the mechanistic and reductionist view.
-That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe, obviously, holds only in principle: in practice, we cannot have such descriptions.
+That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe holds only in principle: obviously, in practice, we cannot have such descriptions.
 However, this contrast between the quantum and the classical view might pinpoint a direction of the research.
-At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system in the longer term with a classical character in the shorter term.
-This might even be a further perspective under which to study the unification between quantum and classical mechanics.
+At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system (standing for the evolution of life on Earth) in the longer term with a classical character in the shorter term.
+This might even be a further perspective under which to study the possibility of unifying quantum and classical mechanics.
 2 Time-symmetric explanation of the quantum speedup
@@
 We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could also apply to the evolution of the universe toward the value of life.
-This section expands on some facts that, for compactness, were just mentioned in the Introduction.
-Of course, in the universe we are in, the possible teleological character of the evolution of the universe toward the value becomes the possible teleological character of the evolution of the universe toward life.
-Then, it is easily seen that this possible teleological character must be inherited by the evolution of life on Earth.
-In principle, the latter evolution can be derived from the evolution of the universe toward life by tracing all the rest.
+This section expands on some facts that, for compactness, were just mentioned, or omitted, in the Introduction.
+Let us show that the possible teleological character of the evolution of the universe toward life must be inherited by the evolution of life on Earth.
+In principle, the latter evolution can be derived from that of the universe toward life by tracing out all the rest.
 Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life.
@@
 Then, just for consistency, the evolutions of species would also be the same.
-Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; it will be very destabilizing with respect to the present way of viewing the evolutions of the living.
+Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; of course, it will be very destabilizing with respect to the present way of viewing both the evolutions of the living and physical reality.
 Let us begin with the evolution of species.
 It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs who did not fly.
-Very light, strong but flexible, cave and air-proof, even at that time feathers were ideal for flight, very close to those of the future birds.
-So why anticipate something designed for flight many millions of years in advance?
+Very light, strong but flexible, cave and air-proof, from the fossils found, even at that time feathers were ideal for flight, very close to those of the future birds.
+So, why anticipate something designed for flight many millions of years in advance?
 The usual explanation, of course, is causal.
@@
 All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred.
+In our possibly biased opinion, this mutual causality explanation of the anticipation in question would be much more convincing than the customary causal one.
 See our preprint [6] for a more comprehensive discussion of the teleological notion.
-All this is obviously destabilizing; it is a way of viewing the time-development of macroscopic reality that goes against the current scientific (causal) way of viewing it.
-Indeed, it revamps the natural philosophy views that were ascribed by science to the category of the unscientific beliefs.
-To the above example of teleological evolution of life, which concerned the evolution of the DNA, we should add another example that concerns the environment of the evolution of life.
-For example, what we humans are today has been affected by the drying up of an area of Africa in which our ancestors lived on trees three million years ago.
-Of course, there is nothing destabilizing in this.
-However, the time-symmetric fact that what we humans are today is also affected by what will concern our descendants in three million years’ time might be very difficult to digest.
-Indeed, it would profoundly change our (causal) way of viewing how things evolve in time.
-Of course, all this is under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle and is thus hypothetical.
-What we would consider certain is the teleological character of quantum algorithms.
-Things could be further explored by investigating, either theoretically or by means of toy models, the character of the “tracing over all the rest”.
-We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing over all the rest.
-We should look for a way of obtaining an evolution of the subsystem that is teleological in the longer term and (about) classical in the shorter term.
-If this were possible, of course, it would go in favor of the above quantum cosmological reading of that version of the Anthropic Principle.
-By the way, it might also be a way of addressing the problem of unifying quantum and classical theory.
-Another question is how to describe a macroscopic evolution endowed with quantum speedup.
+Another possible consequence is for the fact that, according to the present theory, a teleological evolution is an evolution endowed with quantum speedup.
+This could contribute to explaining why some steps of the evolutions of the living have been so rapid to defy Darwinian gradualism, see [18].
+Have been of this kind, in particular, the Cambrian Explosion and modern ecological adaptation.
+Let us also recall what, here, is a very interesting observation of the French biologist Rémy Chauvin [19].
+He pointed out the possibility that some strangely fast evolutionary changes of ants could happen before their being evolutionary advantages.
+As is well known, this biologist merged important contributions to mainstream science with unconventional positions much criticized by colleagues.
+In addition to highlighting the possibility that some evolutionary changes are teleological in character, another unconventional position was his public interest in parapsychology (he wrote books on the subject).
+Note the parallel with the present situation.
+Let us consider the advance knowledge, on the part of the problem-solver, of any of the possible halves of the information about the solution of the problem she will read in the future.
+Besides being a completely technical notion at the basis of the present teleological notion (the problem-solver is a logical agent), under the quantum cosmological assumption it could well be the physical basis of the parapsychological notion of precognition.
+Of course, being under the quantum cosmological assumption, all the above is hypothetical.
+This is to decouple it from the teleological character of quantum algorithms, which we would consider certain.
+Note that a teleological character of the evolutions of the living on the Earth could be further explored by investigating, either theoretically or by means of toy models, the character of the operation of “tracing out all the rest”.
+We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the almost autonomous evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing out all the rest.
+We should look for a way of obtaining an evolution of the subsystem that is teleological (thus quantum) in the longer term and (about) classical in the shorter term.
+If this were possible, it would go in favor of the quantum cosmological reading of the Anthropic Principle we are dealing with.
+By the way, this could also be another perspective under which to address the problem of unifying quantum and classical theory.
+Another question is how to describe a macroscopic evolution of life endowed with quantum speedup.
 Of course, classical physics would turn out to be unable to describe not only the processes of the microscopic world but also the teleological evolutions of the macroscopic world.
@@
 According to the present theory, this discovery would be even more fundamental than currently thought.
-The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [19] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of an attempt to unify relativity and quantum theory.
+The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [21] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of his well known attempt to unify relativity and quantum theory.
 Then, there was the introduction of the notion of logically reversible computation by Bennett [22] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions.
@@
 Let us distinguish between what should be certain and what should be plausible.
-About what should be the certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics.
+What should be certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics.
 (ii) It provides a theoretical confirmation of the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it.
 (iii) It highlights the teleological character of quantum algorithms; given the negative view that the teleological notion has in today’s science and the scope of this notion, this should be the most important result.
-About what is just plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
+What should be plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
+In conclusion, the present work should show that there is another possible view of physical reality out there, outside the current view.
+We note that this might not only be a curiosity.
+A revamping of the abandoned notion of the relation between evolution and functional structure of the living might have a profound impact on today’s evolutionary biology and biology.
 Appendix

### v7

@@
 Abstract
-We show that the quantum computational speedup is due to the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
+We show that the quantum computational speedup of quantum algorithms is due to their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
 We also show that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living for which the teleological notion was originally conceived.
 1 Introduction
-Let us briefly retrace the two and a half millennia history of the teleological notion: attraction exercised by a final cause.
-Plato and Aristotle applied it the functional structure of the living (Aristotle: teeth were created for the purpose of chewing); of course, at the time, there was no notion of natural evolution.
-With the advent of this notion, it was applied to the evolutions of the living by a school of natural philosophy.
-It was eventually dismissed by science (it was classical science) for the alleged absence of a physical basis, even ascribed, for its finalistic character, to the category of beliefs.
-All this was before the advent of quantum mechanics, but has remained almost unchanged with it.
-Today, the situation of the teleological notion is somewhat complex.
-However, to the best of our knowledge, outside the present work, in physics, there are no researches on it.
-Let us round things up by saying that in both physics and the natural sciences there is a de facto disbelief of the retrocasusality and finalism implied in the notion.
-In this work, we shall show that the relatively recent discovery of the quantum computational speedup should change this de facto position of science.
-This is because quantum algorithms turn out to be teleological evolutions: indeed, evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
-Since showing this will be simple, the fact that, forty years after the discovery of the quantum speedup, it has never been noticed before could be ascribed to that de facto position of science.
-We, with coauthors, already went against this position in the evolutionary, retrocausal, approach to the quantum computational speedup [1, 2, 3, 4, 5, 6].
-In it, see [4], we showed that all the quantum algorithms examined, comprising the major ones, obeyed the following rule:
+Once applied to the evolutions of the living by natural philosophy, the notion of teleological evolution—of an evolution attracted by a final goal—was dismissed with the advent of modern science for the alleged lack of a physical basis.
+For its finalistic character, it was also ascribed to the category of beliefs.
+In this work, we shall show that the relatively recent discovery of the quantum computational speedup, the fact that quantum algorithms can solve problems with fewer logical operations (computation steps) than the minimum classically possible, should change this position of science.
+Indeed, we shall show that the quantum speedup of quantum algorithms is due to their teleological character.
+In previous works of coauthors and ours [1, 2, 3, 4, 5, 6], we developed an unconventional, time-symmetric approach to the quantum speedup. In particular, in [4], we showed that:
+In quantum algorithms, it is as if the problem-solver, before beginning her problem-solving action, knew in advance any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to compute the solution with fewer computation steps; the number of computation steps needed to solve the problem in an optimal quantum way is that needed to solve it in an optimal logical way benefiting from the advanced knowledge in question.
+We have called the above the advanced knowledge rule.
+In the above publications, it was given as an empirical rule holding for all the quantum algorithms examined, which comprised the major ones.
+In the present work, we should have demonstrated it.
+Note that this rule, in addition to providing a unified quantitative explanation of all quantum speedups that is absent in mainstream quantum information, implies the teleological character of quantum algorithms.
+Indeed, according to it, they are evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce and measure in the future (the solution of the problem again): Without parentheses and their contents, this can well be considered the definition of teleological evolution.
+This also clarifies the finalistic character of the notion.
+A teleological evolution is not attracted by some superior scientifically undefinable goal, as argued by the opponents of the notion, but, quite simply, by its own end product.
+Note that the advanced knowledge rule establishes the number of computation steps needed to solve an (oracle) problem in an optimal quantum way.
+Whether the quantum algorithm is optimal is another thing.
+We can say that the advanced knowledge rule tells us the room for the quantum speedup.
+Of course, the advanced knowledge rule, its necessarily implying a form of retrocausality, is a destabilizing notion. Let us address in the first place this critical aspect of the work.
+Knowing in advance, on the part of the problem-solver, any of the possible halves of the information about the solution she will produce and measure in the future and using this knowledge to produce the solution with fewer computation steps is obviously a causal loop.
+As there is a plurality of ways of taking half of the information about the solution, we can say that, according to the advanced knowledge rule, the quantum speedup is explained by the mutually exclusive or of these causal loops.
+As far as we know, causal loops are completely absent from ordinary non-relativistic quantum mechanics.
+Of course, the current quantum description of quantum algorithms is causal in character—i.e.
+with causality always going in the same time-direction.
+Naturally, also this quantum description explains the quantum speedup of each quantum algorithm, although in a different way for each quantum algorithm—we shall go back to this further below.
+We have to justify the discrepancy between what is described by the advanced knowledge rule and the current quantum description of quantum algorithms.
+It turns out that the current quantum description is incomplete and, once completed, it comprises the advanced knowledge rule.
+It is incomplete in a double way; the first is trivial (but should be amended first), the second is fundamental.
+The trivial incompleteness consists in the fact that the quantum description is limited to the unitary computation of the solution where the quantum speedup resides and the final measurement of the latter.
+Of course, the canonically complete quantum description of a quantum process consists of initial measurement, unitary evolution, and final measurement.
+This incompleteness is also trivially amended.
+To the quantum description of the process of solving the problem, we should add that of the upstream process of setting the problem.
+An initial measurement of the problem-setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and final measurement of the corresponding solution follow [^1] We exemplify what we mean by problem-setting.
+Consider the following problem: Bob, the problem-setter, hides a ball in a chest of drawers, Alice, the problem-solver, should locate it by opening drawers.
+The number of the drawer with the ball selected by Bob is the problem-setting..
+We call this the ordinary quantum description of the quantum algorithm; “ordinary” because it is still the causal quantum description of mainstream science.
+This ordinary quantum description is still incomplete, this time in the fundamental way.
+Further below, we shall see that it is exactly the same incompleteness lamented by Einstein et al in the case of the quantum description of quantum nonlocality [7].
+Before proceeding, let us simplify the situation by requiring that the solution of the problem is an invertible function of the problem-setting and is always determined with certainty (of course, the problem-setting and the solution are two commuting observables).
+We should also keep in mind that the unitary computation of the solution never changes the problem-setting.
+With the above, we can go to the amendment of the fundamental incompleteness of the quantum description we are dealing with.
+It turns out that this is done by time-symmetrizing it.
+It is a time-symmetrization inspired to time-symmetric quantum mechanics [8, 9, 10].
+As we shall see, the time-symmetrization in question also turns out to be mandatory to satisfy an atemporal symmetry implicit in quantum algorithms and highlighted by their present (ordinary) quantum description.
+Things go as follows.
+Ordinarily, the information that specifies the problem-setting and the corresponding solution is all selected by the initial measurement of the problem-setting.
+We should instead require that this selection evenly shares between the initial measurement of the problem-setting and the final measurement of the solution.
+As there is a plurality of ways of evenly sharing, we should take their quantum superposition.
+In each time-symmetrization instance, the outcome of the initial measurement should propagate forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement.
+Time-symmetrically, the outcome of the final measurement should propagate backward in time, by the inverse of the unitary computation of the solution, until replacing the previous outcome of the initial measurement.
+As we shall see, this backward in time propagation tells the problem solver, before she begins her problem-solving action, one of the possible halves of the information about the solution.
+Indeed, the selection performed by the final measurement affects the initial measurement outcome, changing the initial complete ignorance of the solution on the part of the problem-solver into knowledge of one of the possible halves of the information about it [^2] On the possibility that a future selection affects a past measurement outcome see [10]..
+She can then use this knowledge to produce the solution with fewer computation steps.
+In other words, this time-symmetrization generates the quantum superposition of the causal loops whose mutually exclusive or is implicit in the advanced knowledge rule.
+Besides reconciling the ordinary quantum description of the quantum algorithms with the advanced knowledge rule, this is the demonstration of the latter—we should keep in mind that the time-symmetrization we are dealing with is mandatory.
+By the way, we shall provide another, independent, demonstration of the advanced knowledge rule.
+The fact that the ordinary quantum description of the quantum algorithm involves only the measurements of commuting observables (the problem-setting and the solution of the problem) makes it subject to classical logic.
+Now, as we shall see, the classical logic description of the fact that the quantum algorithm computes the solution with fewer computation steps than the minimum possible in the classical case very directly implies the advanced knowledge rule.
+It is interesting to provide also this demonstration because it very clearly highlights the interplay between time-forward and time-backward causality in unitary quantum evolutions that produces the quantum speedup.
+Let us make another specification in view of what will follow.
+Of course, the sending of information backward in time operated by the time-symmetrization of the ordinary quantum description of the quantum algorithm must be along a time-reversible process.
+Here, it is the unitary computation of the solution.
+The special character of its time-reversibility should be underlined.
+Unlike the classical time-revesibility, which is necessarily an idealization, the present kind is both ideal and objective.
+This is, of course, an essentially quantum feature that stands at the basis of the present time-symmetric explanation of the quantum speedup.
+We have seen that demonstrating the advanced knowledge rule also amounts to demonstrating the teleological character of quantum algorithms.
+Of course, this would already falsify the current opinion that teleological evolutions cannot have a physical basis.
+However, their present physical basis is necessarily in the quantum world and the teleological notion was applied to the evolutions of the living.
+We shall show that, under the quantum cosmological assumption and for the Fine-tuned Version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms identically applies to the evolutions of the living.
+Under the quantum cosmological assumption, the evolution of the universe toward life becomes quantum.
+For that version of the Anthropic principle, it acquires the form of a quantum algorithm and, as such, can be endowed with quantum speed up and consequently be teleological in character.
+The evolution of life on Earth, in principle the trace over all the rest of the former evolution, and, with it, the evolutions of life, would inherit that possible teleological character.
+In summary, the time-symmetrization we are dealing with (and the corresponding completion of the quantum description) turns out to have a special unification power.
+Besides providing a unified quantitative explanation of all quantum speedups (the one provided by the advanced knowledge rule), it unifies the explanation of the quantum speedups with that of quantum nonlocality and, under the quantum cosmological assumption, it unifies the teleological character of quantum algorithms with that of the (macroscopic) evolutions of the living for which the teleological notion was originally conceived.
+2 Time-symmetric explanation of the quantum speedup
+We provide a description of quantum algorithms and their quantum speedups open to the time-symmetric coexistence of causality and retrocausality along unitary quantum evolutions.
+In this section, we provide it at a conceptual level, but making reference to the formalizations postponed in an Appendix.
+The section is divided into subsections corresponding to the various arguments.
+2.1 Critique of the current way of viewing quantum computation
+Let us first acknowledge that that of the quantum computational speedup has been one of the fundamental discoveries of the past century.
+Together with that of quantum communication, it has created an entirely new branch of physics.
+However, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation, likely related to the rejection of the notion of retrocausality.
+We repeat here for convenience the advanced knowledge rule; of course, it plays an essential role in the present work.
 In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
 The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
-We called the rule in question the advanced knowledge rule.
-Note that knowing in advance one of the possible halves of the information about the solution one will produce and measure in the future and benefiting from this knowledge to produce the solution with fewer computation steps is a causal loop.
-We can therefore say that, according to this rule, the quantum speedup is explained by a mutually exclusive or of causal loops, one for each possible way of taking half of the information about the solution.
-By the way, here and in the following, by quantum speedup we will always mean its certain form, the fact that the quantum algorithm needs fewer computation steps than classically possible.
-Of course, this is not the case of most quantum algorithms, which need fewer computation steps than their best classical counterpart, but whether the latter is the best classically possible is not known.
-We should also note that the teleological character of quantum algorithms—their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again)—is implicit in the rule in question.
-Also the notion of finalism acquires a precise definition.
-The teleological evolution would not be attracted by some scientifically undefinable goal but, very simply, by its own end product (in quantum algorithms, the solution of the problem).
-We will see further on that the physical definition of teleological evolution must be that of an evolution between an initial and a final measurement outcome.
-Until now, this advanced knowledge rule was, so to speak, an empirical rule shown to be satisfied by all major quantum algorithms.
-In the present work, we should have demonstrated it, incidentally, in a surprisingly simple way.
-Of course, we would have consequently demonstrated the teleological character of quantum algorithms.
-As a matter of fact, we will provide two redundant demonstrations of this rule, which should not be despised given its destabilizing character. Let us outline them.
-One is an application of classical logic to what the quantum algorithms does: providing a quantum speedup with respect to what is classically possible.
-By the way, we will have to use a quantum description of the quantum algorithms that is not the customary one.
-The latter is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution.
-It is therefore incomplete; the canonically complete description of a quantum process consists of initial measurement, unitary evolution, and final measurement.
-We complete this customary quantum description by adding, to the description of the process of solving the problem, that of the upstream process of setting the problem.
-The result is: an initial measurement of the problem setting in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random; the unitary computation and measurement of the corresponding solution follow.
-Note that this quantum description only involves the measurements of two commuting observables (the problem-setting and the solution of the problem); therefore, it should be subject to classical logic.
-Having said this, the outline of the first demonstration can be very short.
-As we shall see in the next section, under the present more complete quantum description of the quantum algorithm, the classical logic description of the quantum speedup turns out to be indeed the advanced knowledge rule.
-Thus, in particular, this classical logic description of the quantum algorithm would tell more than its present quantum description; we should conclude that the latter is still incomplete.
-Indeed, it is simply completed by time-symmetrizing it.
-We shall see that the time-symmetrization of the more complete quantum description of the quantum algorithm generates the quantum superposition of the causal loops involved in the advanced knowledge rule.
-The second demonstration of the advanced knowledge rule simply comes from the fact that the time-symmetrization of the more complete quantum description of the quantum algorithm turns out to be mandatory for another reason independent from the one seen so far.
-As we shall see, it is to satisfy an atemporal symmetry implicit in this description.
-As we have seen, this time-symmetrization yields the physical correspondent of the advanced knowledge rule—we have thus a demonstration of the rule.
-Naturally, either demonstration of the advanced knowledge rule is also a demonstration of the teleological character of quantum algorithms, their being evolutions toward a goal (the solution of the problem) with an attractor in the very goal they will produce in the future (the solution of the problem again).
-A further step, important for the possibility of extending the teleological character of quantum algorithms to the evolutions of the living, is the following generalization.
-For simplicity, let us limit ourselves to the case that, in the quantum algorithm, the solution is an invertible function of the problem-setting and is determined with certainty.
-We can then say that the quantum algorithm unitarily builds up one-to-one correlation between two initially unrelated observables—the problem-setting and the solution of the problem—without ever changing the eigenvalue of the observable selected by the initial measurement—the problem-setting.
-We shall show that any building up of quantum correlation of this kind, also outside quantum computation, can occur with quantum speedup and therefore be teleological in character.
-Note that the “can occur” means that there is certainly the room for achieving a quantum speedup, but, of course, whether it is achieved, in full or in part or not at all, depends on the character of the unitary evolution that builds up the quantum correlation.
-By the way, in quantum search algorithms, achieving the quantum speedup is not difficult: [7] shows that even a random quantum search for the solution yields a just suboptimal quantum speedup.
-We go now to the evolutions of the living for which the teleological notion was conceived.
-First, let us note that the dismissal of the teleological notion was not without (scientific) regret; indeed, it meant doing without a possibly enlightening relation between the notion of evolution of life and that of the functional structure of the living.
-We cite two authoritative objections to this dismissal.
-One is a famous sentence by the evolutionary biologist Haldane [8]:
-Teleology is like mistress to a biologist, he cannot live without her, but is unwilling to be seen with her in public.
-The other is the position of the natural philosopher Nagel [9], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view.
-In our possibly biased opinion, this would be a case in which common sense prevails over a de facto no-retrocausality precept that, according to the present work, would be a prejudice.
-Let us now go to what the teleological character of quantum algorithms can say to the evolutions of the living.
-Naturally, the first observation must be on the enormous difference between quantum algorithms and the evolutions of the living.
-The former have maximum durations of a few dozen microseconds and a number of quantum bits limited to the order of one thousand; beyond that, in today’s implementations, there is quantum decoherence.
-The latter have a duration of billions of years and can have the size of the Earth.
-However, the very context we are dealing with suggests a way out.
-The difference is enormous in the usual classical way of viewing the macroscopic reality, but disappears under the quantum cosmological assumption that the universe evolves in a unitary quantum way [10, 11].
-Furthermore, surprisingly, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle [11], the evolution of the universe toward the “value of life” (in particular“existence of sentient life” and “absence of life”) could acquire the form of a quantum algorithm and, as such, it could occur with quantum speedup and thus be teleological in character.
-Let us describe how things should go.
-The Fine-tuned universe version of the Anthropic Principle is about the extraordinary theoretical fact that the least change of the value that the fundamental physical constants have in the universe we are in would have developed a relatively trivial universe incapable of producing life.
-Under the quantum cosmological assumption, this version of the Principle can be formalized as follows.
-We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life”; the latter should assume in particular the two values: “existence of sentient life” and “absence of life”.
-Of course, at the beginning of time, these two observables must be unrelated: no matter the value of the fundamental constant, the value of life is “absence of life”.
-At the time of today, the unitary evolution of the universe toward the value of life has built up one-to-one quantum correlation between the value of the fundamental constants since the beginning of time and the value of life, of course, without ever changing the former value.
-We can therefore say that the evolution of the universe toward the value of life can occur with quantum speedup and thus be teleological in character; the “can occur” is in the sense that there is certainly the room for its occurring with quantum speedup.
-Of course, whether the quantum speedup is achieved depends on the character of the unitary evolution.
-For what was said about quantum random search algorithms, we might think that, if the search for the value of life by the unitary evolution of the universe is random, it should occur with a quantum speedup.
-In the universe we are in, the evolution of the universe toward the value of life becomes evolution of the universe toward life; of course, all the above also applies to the latter evolution.
-We shall see that, as a consequence, the evolution of life on Earth, in principle the trace over all the rest of the evolution of the universe toward life, would inherit the possible teleological character of the evolution of the universe toward life.
-For consistency, the same would apply to the evolution of species.
-Note that the fact that a teleological evolution is also an evolution endowed with a quantum speedup is a novelty for the teleological notion.
-It could contribute to explaining steps of natural evolutions that are too rapid to be explained by the progressive character of Darwinian evolutions.
-Summing up, while the teleological character of quantum algorithms, as we shall see, should be certain, the fact that—under the quantum cosmological assumption—it would extend to the evolutions of the living should be considered, at most, as plausible as the quantum cosmological assumption.
-Anyway, all this should be quite different from the present widespread belief that the notion of a teleological evolution of the living is an absurdity.
-At least, it should cast a doubt on this position of science.
-Let us eventually note that, so far, holism and finalism have been the enemy of the reductionist and mechanistic vision of science.
-In the present theory of the quantum computational speedup, they would acquire a physical basis.
-Of course, this would not touch on the immense practical value of the mechanistic and reductionist view.
-That the description of the evolution of any system is the trace over all the rest of the quantum description of the evolution of the universe holds only in principle: obviously, in practice, we cannot have such descriptions.
-However, this contrast between the quantum and the classical view might pinpoint a direction of the research.
-At a toy model, or theoretical, level, it could be interesting to study how this “tracing over all the rest” conciliates a teleological character of the evolution of a system (standing for the evolution of life on Earth) in the longer term with a classical character in the shorter term.
-This might even be a further perspective under which to study the possibility of unifying quantum and classical mechanics.
-2 Time-symmetric explanation of the quantum speedup
-We provide a description of quantum algorithms and their quantum speedups open to the coexistence of causality and retrocausality.
-Since this is a way of thinking quantum computation that is radically different from the usual one, we provide it first at a conceptual level.
-All formalizations are postponed in the Appendix.
-The section is divided into subsections corresponding to the various arguments.
-Let us first acknowledge the fact that that of the quantum computational speedup has been one of the fundamental discoveries of the past century.
-Together with that of quantum communication, it has created an entirely new branch of physics.
-However, for the sake of clarity, we will be completely frank in criticizing what we think are the shortcomings of the current way of viewing quantum computation.
-2.1 On the explanation of the quantum computational speedup
-We repeat here for convenience the advanced knowledge rule.
-In quantum algorithms, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, any of the possible halves of the information that specifies the solution of the problem she will produce and measure in the future and could use this knowledge to produce the solution with fewer computation steps.
-The number of computation steps needed in the optimal quantum case is that needed in the optimal logical case benefiting from the advanced knowledge in question.
-Of course, this rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information.
+This rule would offer a unified quantitative explanation of the quantum speedups that is still lacking in mainstream quantum information.
 Let us discuss this point.
-To the best of our knowledge, the absence of a unified quantitative explanation has never been acknowledged by the scientific community.
+To the best of our knowledge, this absence has never been acknowledged by the scientific community.
 People currently say that there is an explanation of the quantum speedup; it would be that of quantum parallel computation, the fact that many computations are performed at the same time in quantum superposition.
-Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, cannot be considered its explanation.
-By the way, what distinguishes science from opinion is, in particular, its quantitative character.
-The fact that the most popular explanation of the quantum speedup, a quantitative feature, is qualitative in character, would seem to be rather unscientific.
-The only justification we can see is that this is a way of removing the fact that, forty years after its discovery—an explanation of the quantum speedup is still missing.
-This is naturally our opinion.
-However, as a matter of fact, outside our (unconventional) retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
-Of course, as already noted, if the only way of providing the unified quantitative explanation in question is by resorting to retrocausality and finalism, the de facto rejection of these notions by the scientific community might have played a role.
-Neither can be considered an explanation the fact that each quantum algorithm naturally explains its quantum speedup.
+Indeed, this captures the essential character of quantum computation, but, being a qualitative explanation of the quantum speedup, which is a quantitative feature, it cannot be considered its explanation.
+Of course, science essentially differs from opinion because of its quantitative character; then, settling for a qualitative explanation of a quantitative feature is rather unusual.
+This is our opinion, of course.
+However, as a matter of fact, outside the present retrocasusal approach to the quantum speedup, in the literature, there is no trace whatsoever of any search for a unifying-quantitative explanation of it.
+Neither can be considered a satisfactory explanation of the quantum speedup the fact that each quantum algorithm naturally explains its quantum speedup.
 In fact, it would be an explanation completely different for different quantum algorithms.
 It would be like having different explanations for different throws of a stone without knowing that they all describe a parabola.
-Obviously, unification is a fundamental objective of any description of physical phenomena.
-Summing up, in our judgment, for what concerns the current mainstream explanation of the quantum speedup, it is the case of saying that the emperor is naked.
+Naturally, unification of explanations is a fundamental objective of science.
 Of course, the advanced knowledge rule would exactly be that unifying-quantitative explanation.
-However, to date, it has had a twofold problem: first, although shown to apply to all the major quantum algorithms, it has not been demonstrated; second, it goes against the de facto ban on retrocausality.
-In this work, we should have demonstrated it, moreover, in a surprisingly simple way.
-As a matter of fact, we shall be dealing with two independent demonstrations, which is not to be despised because of the destabilizing character of the rule.
-2.2 First demonstration of the advanced knoledge rule
-We have already provided in the Introduction the quantum description of quantum algorithms extended to the process of solving the problem.
+In this section, we shall provide the two demonstrations of it, beginning with the one relying on the classical logic description of the quantum speedup.
+2.2 First demonstration of the advanced knowledge rule
+Let us recall the quantum description of quantum algorithms extended to the process of solving the problem already provided in the Introduction:
 An initial measurement in a quantum superposition (or, indifferently, mixture) of all the possible problem-settings selects one of them at random [^3] The problem-setter could unitarily change it at will; this, here, would change nothing..
@@
 For the sake of simplicity, we consider the case that the solution is an invertible function of the problem-setting and is determined with certainty.
-Now, we should note that this quantum description of the quantum algorithm is with respect to the customary observer external to the quantum process.
+Table I of Section (1) of the Appendix provides the formalization of this extended quantum description in the simplest instance of Grover algorithm.
+Here, the problem is as follows: Bob, the problem-setter, hides a ball in a chest of four drawers, Alice, the problem-solver, is to locate it (i.e.
+identify the number of the drawer with the ball) by opening drawers.
+Note that the quantum description of the quantum algorithm of Table I is with respect to the customary observer external to the quantum process.
 To present ends, we should relativize it with respect to the problem-solver who, for the fact of computing the solution of the problem, is part of the process.
-Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the completed quantum description of the quantum algorithm with respect to her.
-Now, from her, the problem-setting selected by the initial measurement should be concealed: as we shall see in the following, it would tell her the solution of the problem before she begins her problem-solving action.
+Of course, the advanced knowledge of one of the possible halves of the information about the solution we are dealing with will have to appear in the (completed) quantum description of the quantum algorithm with respect to her.
+Now, the problem-setting selected by the initial measurement should be concealed from her: as we shall see, it would tell her the solution of the problem before she begins her problem-solving action: Of course, at this time, instead, she is completely ignorant of the solution.
 Since the unitary computation of the solution never changes the problem-setting selected by the initial measurement, this concealment is easy.
-It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of her unitary computation of the solution; see Section (2) of the Appendix.
+It suffices to postpone the projection of the quantum state associated with the initial measurement to the end of her unitary computation of the solution; see Table II of Section (2) of the Appendix.
 Let us call this extended quantum description with respect to the problem-solver “the ordinary quantum description of the quantum algorithm”.
 It is ordinary because it is still the customary “causal” quantum description of mainstream quantum mechanics—where causality always goes in the same time-direction.
-We can go now to the first demonstration of the advanced knowledge rule.
-It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that is less than the minimum possible in the classical case.
-By the way, involving only the measurements of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
-To present ends, we need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible.
+We can now go to the first demonstration of the advanced knowledge rule.
+It simply consists in describing in a classical logic way what the quantum algorithm does—computing the solution of the problem with a number of computation steps that are less than the minimum possible in the classical case.
+By the way, since their initial and final measurements are of commuting observables (the problem-setting and the solution of the problem) quantum algorithms, and what they do, should be subject to classical logic.
+Now, we have to choose the quantum algorithm to work with.
+We need a quantum algorithm that certainly gives a quantum speedup with respect to what is classically possible.
 Now, most quantum algorithms yield a quantum speedup with respect to their best classical counterpart, but whether the latter is the best possible is not known.
 Excluding elementary one-computation-step quantum algorithms whose dimensions do not scale up, we find our choice limited to Grover’s quantum search algorithm [13].
-Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) It is a most universal quantum algorithm that can solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup.
-(ii) In Long’s [13] version, it always provides the solution with certainty; this simplifies things.
-(iii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization.
-(iv) It is optimal in character; this, of course, constitutes an important reference.
-By the way, note that the notion of searching for something and finding it is most universal and that this universality is completely represented by Grover algorithm.
-We show that the classical logic description of the quantum speedup delivered by Grover algorithm very directly implies the advanced knowledge rule.
-We begin with the simplest instance of Grover algorithm.
-As we shall soon see, everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
+Besides yielding a well determined quantum speedup with respect to what is classically possible, it has the following convenient features: (i) In Long’s version [14], it always provides the solution with certainty; this simplifies things.
+(ii) In it, the solution is an invertible function of the problem-setting; this simplifies its time-symmetrization.
+(iii) It is optimal in character; this, of course, constitutes an important reference.
+(iv) It is a most universal quantum algorithm: it can easily be adapted to solve any NP problem with a quadratic quantum speedup; more than a single quantum algorithm, it is an entire test bed on which to test the present theory of the quantum speedup.
+Let us show that (under classical logic) the quantum speedup over what is classically possible delivered by Grover algorithm logically implies the advanced knowledge rule.
+We begin with the simplest instance of this quantum algorithm: everything immediately generalizes first to the full fledged Grover algorithm and then to most general forms of quantum correlation (in quantum algorithms, that between the problem-setting and the solution of the problem).
 The problem to be solved is as follows.
@@
 Alice, the problem-solver, is to locate the ball by opening drawers.
-In the worst classical case, she has to open three drawers.
-If the ball is only in the third drawer opened, she has solved the problem.
+In the worst classical case, she has to open three drawers: If the ball is only in the third drawer opened, she has solved the problem.
 If not, it must be in the only drawer not yet opened, and she has solved the problem of locating the ball as well.
@@
 Obviously, always solving the problem by opening just one drawer would be impossible in the classical case: indeed, we are dealing with a quantum computational speedup.
-There is a classical logic equivalent to the fact that the quantum algorithm always solves the problem by opening a single drawer.
+There is a classical logic equivalent of the fact that the quantum algorithm always solves the problem by opening a single drawer.
 By the way, since the fact that a single drawer is opened is logically true in itself, we can forget the specification that this is done for a quantum superposition of the four drawer numbers.
 The logical equivalence is as follows.
-Logically, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers (one of which, of course, the one with the ball) and took advantage of this knowledge to solve the problem by opening just one of them.
+Logically, it is as if the problem-solver always knew in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers (in a mutually exclusive way one of the three pairs of drawers in which the ball must be) and took advantage of this knowledge to solve the problem by opening just one of them.
 In fact, if the ball is in the opened drawer, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer.
-This also means knowing in advance half of the information that specifies the solution of the problem that will be produced and measured in the future.
+By the way, this also means knowing in advance half of the information that specifies the solution of the problem that will be produced and measured in the future.
 Let us open a parentheses to observe that we are not talking about the quantum algorithm as it were a classical algorithm.
 In other words, here, classical logic should not be confused with classical physics, which might be a temptation in the present case.
-Always knowing in advance that the ball is in a given pair of drawers, in a mutually exclusive way, one of the possible pairs that host the drawer with the ball, is logically necessary to always solve the problem by opening just one drawer.
+Always knowing in advance that the ball is in a given pair of drawers is logically necessary to always solve the problem by opening just one drawer.
 Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution?
@@
 Of course, this is the advanced knowledge rule for the present instance of Grover algorithm, logically implied by the sheer existence of the quantum speedup.
-In summary, the classical logic description of the quantum speedup delivered by (the simplest instance of) Grover algorithm implies the advanced knowledge rule.
 This explanation of the quantum speedup immediately generalizes to any number $N$ of drawers.
@@
 This is, of course, the advanced knowledge rule for the full fledged Grover algorithm.
-Note that, in equivalent terms, the computational complexity of the problem to be solved reduces, from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers.
+Note that, in equivalent terms, the computational complexity of the problem to be solved reduces from searching the ball in $N$ drawers to searching it in $\sqrt{N}$ drawers.
 In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place reduces from $N$ to $\sqrt{N}$.
 Holding for Grover algorithm, the above retrocausal explanation of the quantum speedup must hold for the process of solving with a quadratic quantum speedup any NP problem.
-Here we do not care about the fact that most NP problems are classically solved with a number of computation steps exponential in their size, so that a quadratic quantum speedup can be seen as a little advantage.
-Of course, for the present ends, we are just interested in having a quantum speedup with respect to what is classically possible.
 We shall see that there are further generalizations in the following of this section.
 In view of them, note that, in order to derive the present explanation of the quantum speedup, we never needed to know how Grover algorithm achieves the latter, we only needed to know that it achieves it.
-2.3 Completing the ordinary quantum description of the quantum algorithms by time-symmetrizing it
+By the way, let us underline the fact that the advanced knowledge rule tells us what can be the quantum speedup in the optimal quantum case.
+Of course, whether this quantum speedup is achieved, or achieved in part, or not at all, depends on the character of the unitary evolution that computes the solution.
+Therefore, we can say that the advanced knowledge rule tells us the room for the quantum speedup.
+Incidentally, achieving a quantum speedup may not be difficult: [15] shows that even a random quantum search for the solution yields just a suboptimal quantum speedup.
+2.3 Completing the ordinary quantum description of quantum algorithms by time-symmetrizing it
 The above subsection also says that the ordinary quantum description of quantum algorithms, which obviously does not describe the advanced knowledge rule, is incomplete.
 Of course, to bring it in line with their classical logic description, it should be completed.
-It turns out that this is simply done by time-symmetrizing it.
+It turns out that this is simply done by time-symmetrizing it; we shall also see that this time-symmetrization is mandatory to satisfy an atemporal symmetry implicit in this quantum description.
 We describe this operation.
 For the sake of simplicity, we consider the case that the solution is determined with certainty and is an invertible function of the problem-setting—exactly as in Grover algorithm.
-We should evenly share between the initial measurement of the problem-setting and the final measurement of the solution the selection of the information that specifies both; of course, in the ordinary quantum description of the quantum algorithm, all the information is selected by the initial measurement.
-Naturally, there is a plurality of possible ways of evenly sharing: we should take their quantum superposition.
+Naturally, in the ordinary quantum description of the quantum algorithm, the initial measurement of the problem-setting selects the information that specifies both the problem-setting and the corresponding solution.
+To time-symmetrize this ordinary quantum description, we should evenly share the selection of this information between the initial measurement of the problem-setting and the final measurement of the solution.
+Since there is a plurality of possible ways of evenly sharing, we should take their quantum superposition.
 In each time-symmetrization instance, the selection performed by the initial measurement propagates forward in time by the unitary computation of the solution until it becomes the state immediately before the final measurement.
-Time-symmetrically, that performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement.
-This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum algorithm.
-As we can see in Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum algorithm it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; she can then use this knowledge to compute the solution with fewer computation steps.
-This is the physical correspondent of one of the logical causal loops whose mutually exclusive or is logically implied by the advanced knowledge rule.
+Time-symmetrically, the selection performed by the final measurement propagates backward in time by the inverse of the unitary computation of the solution until replacing the previous outcome of the initial measurement.
+This backward in time propagation, which inherits both selections, is an instance of the time-symmetrized quantum description.
+As we can see in Table III of Section (3) of the Appendix, in the initial measurement outcome of the time-symmetrized instance of the quantum description, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future (left of the bottom line of Table III diagram); she can then use this knowledge to compute the solution with fewer computation steps (bottom line of Table III diagram or Table IV diagram).
+The full Table III diagram is the physical correspondent of one of the causal loops whose mutually exclusive or is logically implied by the sheer existence of the quantum speedup.
 Eventually, we should take the quantum superposition of all the time-symmetrization instances.
 The time-symmetrized quantum description of the quantum algorithm (i.e.
-its completed quantum description) becomes the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the quantum speedup: So completed, also the quantum description of the quantum algorithm describes the advanced knowledge rule described by its classical logic description—it describes its physical counterpart.
+its completed quantum description) becomes the quantum superposition of the causal loops whose mutually exclusive or is logically implied by the quantum speedup: Thus completed, also the quantum description of the quantum algorithm describes the advanced knowledge rule described by its classical logic description—more precisely it describes its physical counterpart.
 2.4 Second demonstration of the advanced knowledge rule
@@
 In the ordinary quantum description of the quantum algorithm, of course, the unitary computation of the solution never changes the corresponding problem-setting.
-Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it is simultaneous with the final measurement of the solution.
-In this case, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements.
+Therefore, one should be free to postpone the initial measurement of the problem-setting along the subsequent unitary computation of the solution until it becomes simultaneous with the final measurement of the solution.
+At this point, we find ourselves with the simultaneous measurements of two maximally entangled observables (the problem-setting and the solution, one an invertible function of the other), and the time-symmetrization we are dealing with must already be in place to represent the perfect (atemporal) symmetry between these two measurements.
 Of course, there would be no reason why a measurement selected more information than the other.
@@
 This is indeed the second demonstration of the advanced knowledge rule.
-By the way, this result is reminiscent of Wheeler and Feynman’s [14] time-symmetric absorber theory; we would say: with the advantage that the time-symmetrization we are presently dealing with relies on a reversibility of the quantum process that, unlike the classical one, is such in both an ideal and objective way.
-2.5 Origin of the retrocausality
+By the way, this result is reminiscent of Wheeler and Feynman’s [16] time-symmetric absorber theory; we would say: with the advantage that the time-symmetrization we are presently dealing with relies on the reversibility of unitary quantum evolutions, which, unlike classical reversibility, is such in both an ideal and objective way.
+2.5 Origin of the mutual causality
 A quantum algorithm is a relatively complex process.
-A natural question is whether the retrocausality implicit in the classical logic description of its overall result, i.e.
+A natural question is whether the mutual causality implicit in the classical logic description of its overall result, i.e.
 the quantum speedup, derives from some more elementary part of it.
@@
 Since the latter have a classical analog, the essentially quantum generation of a quantum state superposition is the suspect.
-Let us first provide a very intuitive exposition of the fact that the classical logic description of a state superposition necessarily involves the coexistence of causality and retrocausality.
-We resort to a prescientific notion of state superposition.
-It is the legend of the monk seen by his disciples on the two banks of a river at the same time: we should provide a classical logic description of this.
-Being on one and the other bank is possible in two different times but, at the same time, from the ordinary classical logic standpoint, they should be two mutually exclusive possibilities.
-Until now, being on the two banks at the same time would seem to be a logical impossibility.
-However, let us note that it would be impossible in a classical logic description closed to retrocausality.
-Opening classical logic to the coexistence of causality and retrocausality completely changes the situation [^2] By the way, classical logic is neutral for what concerns the notion of causality.
-Associating it with a causality that always goes in the same time-direction, as customary, must be consequent to the de facto no-retrocausality precept.
-Classical logic should also go along with the coexistence of causality and retrocausality..
-Indeed, the monk could have gone to one of the two banks along a certain path, then (then in causal order) he could have gone backward in time undoing the path just followed (like in a reversed film); then, going forward in time again, he could have reached the other bank along a different path.
-This causality zigzag, of course, makes its being on the two banks at the same time logically possible.
-Let us now bring the above classical logic description of a state superposition to a quantum state superposition.
+Let us show that the classical logic description of the unitary generation of a quantum state superposition necessarily involves retrocausality.
 Think of a photon going in quantum superposition through the two branches of a Mach-Zehnder interferometer.
 The two branches should be tuned in such a way that the photon always ends up in the same detector, so that we have a (of course, time-reversible) unitary evolution that goes from before the photon enters the interferometer to immediately before its detection.
-The classical logic description of this can be as follows.
+The classical logic description of this quantum superposition can be as follows.
 In entering the interferometer, the photon takes, say, the upper branch until it reaches the state immediately before detection (by the detector in question).
@@
 Then (in causal order), the photon goes backward in time undoing the path just followed until before entering the interferometer.
-Then, it goes forward in time again taking the lower branch until reaching the detector in question and, this time, being detected.
+Then, it goes forward in time again taking the lower branch until reaching the detector in question, and this time, be detected.
 The zigzag path followed in this way by the photon through the interferometer, with the associated photon phases, of course, describes in a classical logic way the fact that the photon goes through the interferometer in a quantum superposition of the upper and lower branches, eventually ending up in that detector.
-It should be clear that, also in this case, the involvement of retrocausality in the classical logic description is necessary.
+It should be clear that the classical logic description of this unitary generation of a quantum state superposition necessarily involves the coexistence of the two time directions of causality along the same unitary process.
 We should think that the classical logic description of the sequence of generations of quantum state superpositions and interferences that constitutes the unitary computation of the solution, in the end, simplifies into a description that implies the advanced knowledge rule.
-Summarizing, at the root of the logical form of retrocausality we are dealing with there should be the fundamental notion of quantum state superposition, the fact that the classical logic description of its generation necessarily involves retrocausality.
+Summarizing, at the root of the form of mutual-causality we are dealing with there would be the fundamental notion of quantum state superposition.
 By the way, note that this is not an idle speculation as it might even seem.
-Indeed, we are dealing with a form of retrocausality with huge consequences, such as providing a physical basis to the notion of teleological evolution.
+Indeed, we are dealing with a form of coexistence of the two time-directions of causality with huge consequences, such as providing a physical basis for the notion of teleological evolution.
 2.6 Generalization
-In the second demonstration of the advanced knowledge rule, we have shown that a unitary evolution between two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement is incomplete and must be completed by time-symmetrizing it.
-Note that we never needed to look inside the unitary evolution, for example $\hat{\mathbf{U}}_{1,2}$ of Tables I through IV of the Appendix.
-Thus, the need for this time-symmetrization does not only hold for quantum algorithms; there is the following generalization:
+In the second demonstration of the advanced knowledge rule, we have seen that the ordinary quantum description of a quantum algorithm whose solution is an invertible function of the problem-setting and is determined with certainty is incomplete and must be completed by time symmetrizing it.
+One can see that the necessity of this time-symmetrization does not depend on the fact that we are dealing with a quantum algorithm; there is this immediate generalization:
 The ordinary quantum description of any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is incomplete and is completed by time-symmetrizing it.
-Let us call it the quantum correlation time-symmetrization rule.
-The case where the one-to-one correlation is built up by the unitary evolution characterizes the quantum search algorithm, where this building up occurs with quadratic quantum speedup.
-However, the possibility of occurring with quadratic quantum speedup holds for any unitary evolution that responds to the rule in question and builds up one-to-one quantum correlation between two initially unrelated observables.
-Indeed, that this process can occur with quadratic quantum speedup is what the time-symmetrized quantum description of the quantum process says—see the Appendix.
+Indeed, as one can see, to this kind of unitary evolution one can apply the argument for which it must be time-symmetrized to satisfy the atemporal symmetry implicit in it.
+Let us call this rule the quantum correlation time-symmetrization rule.
+In the case that the unitary evolution builds up one-to-one correlation between two initially unrelated observables, as in quantum search algorithms, this building up can always occur with quadratic quantum speedup, in the sense that there is always the room for its occurring with quantum speedup.
+This is what the time-symmetrization of its ordinary quantum description says—see Section (3) of the Appendix.
 Of course, in the case where we are not dealing with a quantum algorithm, the notion of advanced knowledge of one of the possible halves of the information about the solution must be replaced by the equivalent notion of quadratic reduction of the dimension of the Hilbert space in which the search for that one-to-one correlation takes place, from $N$ to $\sqrt{N}$.
-The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what would preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality.
-Here, we have a unitary evolution that spatially separates the subsystems of two already maximally entangled observables.
-With the two measurements performed in the same basis and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
-As shown in [5], the time-symmetrization of the ordinary quantum description of this form of quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [15].
+The case in which the one-to-one quantum correlation is not built up by the unitary evolution because it preexists it (more precisely, what should preexist is the maximum entanglement between the observables of the two measurements) is, for example, that of quantum nonlocality with the two measurement apparatuses aligned.
+Here, the unitary evolution spatially separates the subsystems of two already maximally entangled observables.
+With the two measurement apparatuses aligned and the spatial separation process preceding the first measurement, this responds to the quantum correlation time-symmetrization rule.
+As shown in [5], the time-symmetrization of the ordinary quantum description of this quantum nonlocality yields a quantum superposition of causal loops that locally explain it and play the role of the hidden variables envisaged by Einstein et al [7].
 Indeed, these causal loops locally go from one to the other measurement outcome via the time the two subsystems were not yet space separate, where they locally tell the subsystem of the not yet measured observable the state to assume in view of the future measurement.
 Note that this is exactly what Einstein et al required from their hidden variables.
-That was, of course, without resorting to retrocausality, but under a deterministic view of the measurement process equivalent in this case.
-In summary, this application of the quantum correlation time-symmetrization rule both demonstrates the incompleteness of the ordinary quantum description of quantum nonlocality and provides the hidden variables that complete it.
-Of course, this also means that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
-By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al might be destabilizing for the following reason: After highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [16] and Ekert’s quantum cryptography based on it [17], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
-By the way, if the only possible hidden variables are retrocausal in character as the ones we are dealing with, this dismissal can be ascribed to the already mentioned de facto rejection of retrocausality and finalism (the form it has in philosophy, different from the present trivial form).
-2.7 About the teleological character of the evolutions of the living
-We have already seen in the Introduction that, under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms could also apply to the evolution of the universe toward the value of life.
-This section expands on some facts that, for compactness, were just mentioned, or omitted, in the Introduction.
-Let us show that the possible teleological character of the evolution of the universe toward life must be inherited by the evolution of life on Earth.
-In principle, the latter evolution can be derived from that of the universe toward life by tracing out all the rest.
+That was, of course, without resorting to any mutual-causality, but under a deterministic view of the measurement process equivalent in this case.
+Of course, this should also mean that the incompleteness we are presently ascribing to the ordinary quantum description of quantum algorithms is the same that Einstein and others ascribed to the quantum description of quantum nonlocality.
+By the way, the fact that the present explanation of the quantum speedup also demonstrates the incompleteness of the quantum description highlighted by Einstein et al might also be destabilizing.
+Indeed, after highlighting for the first time quantum nonlocality and inspiring fundamental works like Bell’s theorem [17] and Ekert’s quantum cryptography based on it [18], their argument had been dismissed by the physicists’ community because of the alleged absence of a satisfactory way of completing the quantum description by means of local hidden variables.
+In particular, the argument of the incompleteness of the quantum description has been forgotten.
+2.7 The possible teleological character of the evolutions of the living
+We have seen that the possibility of a teleological character of a quantum algorithm, of its occurring with quantum speedup and thus be teleological in character, depends only on the form of its ordinary quantum description: of the unitary evolution between the initial an the final measurement we need only to know the input and the output.
+We do not need to know its inside.
+All this defines the room for the quantum speedup, which is what is of interest here.
+The unitary evolution might even not be a quantum algorithm.
+Let us show how the teleological character of quantum algorithms could be ported to the (macroscopic) evolutions of the living for which the teleological notion was originally conceived.
+The first thing one should observe is that quantum algorithms belong to the microscopic world of quantum phenomena and, as things are now, have durations of a few dozen of microseconds: more than that, there is quantum decoherence.
+The evolutions of life, instead, can have the size of the earth and have duration of billions of years.
+The distance between the two is immense but, it should be noted, it would be so under the customary classical vision of macroscopic reality.
+Instead, under the quantum cosmological assumption that the universe evolves in a unitary quantum way, this distance vanishes.
+By the way, recall the perfect time-reversibility of unitary quantum evolutions.
+Now there is a surprising fact.
+Under the quantum cosmological assumption and for the Fine-tuned Universe version of the Anthropic Principle, the evolution of the universe toward the value of life acquires the form of a quantum search algorithm and, as such, it can occur with quantum speedup and thus be teleological in character, this in the sense that there is certainly the room for that.
+Let us explain.
+That version of the principle is about the extraordinary theoretical fact that the least change of the value (for values) that the fundamental physical constants has in the universe we are in would have produced a relatively trivial universe unable to develop life.
+Under the quantum cosmological assumption, this can be represented as follows.
+We should introduce two commuting observables of the same size: the value of the fundamental constants and the “value of life” ; the latter should assume, in particular, the values “absence of life” and “existence of sentient life”.
+We should then consider a quantum superposition (or, indifferently, mixture) of all the possible universes, each with its own value of the fundamental constants, all obeying the same fundamental laws (but for the specific value of the fundamental constants).
+Of course, at the beginning of time, in this quantum superposition, these two observables must be unrelated: no matter the value of the fundamental constants, the value of life is “absence of life”.
+We should now consider the unitary evolution of this quantum superposition of all possible universes toward the value of life.
+At the time of today, this evolution must have built up maximal entanglement (say, one-to-one correlation) between the value of the fundamental constants since the beginning of time and the value of life, of course, without ever changing the former value.
+By the way, according to cosmology, at the time of today, the absence of life in the universes that cannot produce it would already be definitive.
+We have seen that this kind of building up of quantum correlation can always occur with quantum speedup (in the sense that there is always the room for it) and therefore be teleological in character.
+In the universe we are in, the evolution of the universe toward the value of life becomes evolution of the universe toward life; and, at the time of today, there are the sentient observers who measure the value of life, also selecting the value of the fundamental constants maximally entangled with it.
+By the way, since the latter value is never changed by the unitary evolution of the universe toward life, all is as if it was measured at the beginning of time.
+We could also think that, if the search for the value of life by the evolution of the universe is random, it would occur with quantum speedup.
+Summarizing, the evolution of the universe toward life could occur with quantum speedup and consequently be teleological in character.
+Let us show that this possibility would then be inherited by the evolution of life on Earth and the evolution of species.
+In principle, the evolution of life on Earth can be derived from the evolution of the universe toward life by tracing out all the rest.
 Conversely, we can speak of an image of the evolution of life on Earth in the evolution of the universe toward life.
@@
 In fact, the latter resources comprise the former.
-In summary, under the quantum cosmological assumption, the evolution of life on Earth could occur with quantum speedup (in the sense that there would be the room for this) and consequently be teleological in character; it might actually occur with quantum speedup if the search for life by the evolution of the universe toward life were random.
+In summary, under the quantum cosmological assumption, the evolution of life on Earth could occur with quantum speedup (in the sense that there would always be the room for it) and consequently be teleological in character; it might actually occur with quantum speedup if the search for life by the evolution of the universe toward life were random.
 Then, just for consistency, the evolutions of species would also be the same.
-Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; of course, it will be very destabilizing with respect to the present way of viewing both the evolutions of the living and physical reality.
-Let us begin with the evolution of species.
-It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs who did not fly.
+Not to hide behind a finger, let us exemplify what a teleological character of the evolution of life would imply; of course, it would be very destabilizing with respect to the present way of viewing both the evolutions of the living and physical reality.
+It is known that, for many millions of years, the ancestors of the birds were feathered dinosaurs that did not fly.
 Very light, strong but flexible, cave and air-proof, from the fossils found, even at that time feathers were ideal for flight, very close to those of the future birds.
@@
 The usual explanation, of course, is causal.
-First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually feathers turned out to be what was needed to conquer the air.
-All this could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred.
-In our possibly biased opinion, this mutual causality explanation of the anticipation in question would be much more convincing than the customary causal one.
-See our preprint [6] for a more comprehensive discussion of the teleological notion.
-Another possible consequence is for the fact that, according to the present theory, a teleological evolution is an evolution endowed with quantum speedup.
-This could contribute to explaining why some steps of the evolutions of the living have been so rapid to defy Darwinian gradualism, see [18].
-Have been of this kind, in particular, the Cambrian Explosion and modern ecological adaptation.
-Let us also recall what, here, is a very interesting observation of the French biologist Rémy Chauvin [19].
-He pointed out the possibility that some strangely fast evolutionary changes of ants could happen before their being evolutionary advantages.
-As is well known, this biologist merged important contributions to mainstream science with unconventional positions much criticized by colleagues.
-In addition to highlighting the possibility that some evolutionary changes are teleological in character, another unconventional position was his public interest in parapsychology (he wrote books on the subject).
-Note the parallel with the present situation.
-Let us consider the advance knowledge, on the part of the problem-solver, of any of the possible halves of the information about the solution of the problem she will read in the future.
-Besides being a completely technical notion at the basis of the present teleological notion (the problem-solver is a logical agent), under the quantum cosmological assumption it could well be the physical basis of the parapsychological notion of precognition.
+First, an ancestor of feathers appeared by random DNA mutation, then evolved as a good thermal insulator and perhaps to speed escape from predators; only eventually (by sheer coincidence) feathers turned out to be what was needed to conquer the air.
+All this (without the parentheses and its content) could remain true in the teleological view with just one, but very significant, addition: Those random mutations that would have led in the future to the conquest of the air would have more likely occurred.
+In our possibly biased opinion, this mutual-causality explanation of the anticipation in question would be much more convincing than the customary causal explanation.
+Indeed, the feathers of the fossils of those non-flying dinosaurs are almost identical to those of their flying descendants.
+Our preprint [6] provides a more extensive discussion of the implications of the teleological notion.
+Another consequence of the present theory is that, according to it, a teleological evolution is an evolution endowed with quantum speedup.
+Of course, this is a feature of teleological evolutions completely absent in their historical notion.
+It could contribute to explaining why some steps of the evolutions of the living have been so rapid to defy Darwinian gradualism, see [19].
+As is well known, have been of this kind, in particular, the Cambrian Explosion and modern ecological adaptation.
 Of course, being under the quantum cosmological assumption, all the above is hypothetical.
-This is to decouple it from the teleological character of quantum algorithms, which we would consider certain.
-Note that a teleological character of the evolutions of the living on the Earth could be further explored by investigating, either theoretically or by means of toy models, the character of the operation of “tracing out all the rest”.
-We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it the almost autonomous evolution of a subsystem of it, standing for the evolution of life on Earth, by tracing out all the rest.
+This is to decouple it from the teleological character of quantum algorithms, which we would consider watertight.
+3 Discussion
+First, let us show that the abandonment of the teleological perspective in the natural sciences was and is not without criticism.
+An example is a famous sentence by the evolutionary biologist Haldane [23]:
+Teleology is like mistress to a biologist, he cannot live without her, but is unwilling to be seen with her in public.
+The other is the position of the natural philosopher Nagel [24], who provides a series of very practical arguments for which the mechanistic, neo-Darwinian conception of nature could not account for the fundamental features of the living, and this in favor of a holistic and teleological view.
+It could be said that, in general, natural philosophy is still pro the teleological notion. The following is a quotation from the natural philosophers Allen and Neal [22]:
+Teleological notions are largely considered ineliminable from modern biological sciences, such as evolutionary biology, genetics, medicine, ethology, and psychiatry, because they play an important explanatory role.
+Against the above, there is the position of the physical science and its de facto rejection of the teleological notion.
+By the way, this is also the position officially adopted by the natural sciences.
+In our possibly biased view, natural philosophy would respond better to empirical evidence and be less affected by what we would consider a de facto prejudice of science.
+Note that a teleological character of the evolutions of the living on the Earth could be further explored by investigating, either theoretically or by means of toy models, the character that the operation of “tracing out all the rest” would acquire in the present context.
+We should consider the unitary quantum evolution of a system endowed with quantum speedup, standing for the evolution of the universe toward life, and derive from it, by tracing out all the rest, the almost autonomous evolution of a subsystem of it, standing for the evolution of life on Earth.
 We should look for a way of obtaining an evolution of the subsystem that is teleological (thus quantum) in the longer term and (about) classical in the shorter term.
-If this were possible, it would go in favor of the quantum cosmological reading of the Anthropic Principle we are dealing with.
-By the way, this could also be another perspective under which to address the problem of unifying quantum and classical theory.
+If this were possible, it would go in favor of the quantum cosmological reading of the Anthropic Principle that we have proposed.
+By the way, this could also be another perspective under which to address the problem of unifying quantum and classical theory, in this case under the present form of time-symmetric quantum mechanics.
+The same might also hold for cosmology.
+After all, under the quantum cosmological assumption, also the physical part of the evolution of the universe toward life should be subject to the Fine-tuned Universe version of the Anthropic Principle.
+Also this part could be endowed with quantum speedup and thus be teleological in character.
+It might be interesting to confront this possibility with the present cosmological notions, naturally comprising their open problems.
+We have already noted that the present notion of teleological evolution implies the quantum speedup of the latter.
+This would, of course, be a novelty with respect to the historical notion of teleological evolution.
+It might even better explain than Darwinian gradualism the extraordinary speed with which some evolution steps occurred.
 Another question is how to describe a macroscopic evolution of life endowed with quantum speedup.
@@
 Indeed, if it were completely autonomous, of course, it would be a unitary quantum evolution itself.
-Let us end the present section by recalling what, to our knowledge, are the key steps that led to the discovery of the quantum computational speedup.
-According to the present theory, this discovery would be even more fundamental than currently thought.
-The first step was showing that computation is possible in the quantum world and introducing the fundamental notion of quantum bit, by Finkelstein [21] in 1969; that was also the first time that quantum physics moved to the world of the complex, indeed to a cellular automaton model of the universe—this was in the context of his well known attempt to unify relativity and quantum theory.
-Then, there was the introduction of the notion of logically reversible computation by Bennett [22] in 1973, of course, a prerequisite for the building up of computations in unitary quantum evolutions.
-The discovery of the superior computational efficiency of the quantum world with respect to that of the classical world in 1982 is due to Feynman [23].
-Still in 1982, there has been the introduction of the notion of conservative logic by Fredkin and Toffoli [24]: with it, we have the logically reversible gates needed to perform quantum computations.
-The seminal quantum algorithm that delivers a quantum speedup was conceived by Deutsch [25] in 1985.
-Note the importance of the role that the notion of reversibility played in the development of quantum computation.
-The present work could be put on the same path.
-It could be said that the present time-symmetric form of causality exploits the objectively ideal time-reversibility of unitary quantum evolutions in the deepest possible way.
-3 Precedents
-We have seen that applying the quantum correlation time-symmetrization rule to quantum nonlocality provides a local explanation of it. Now, we recall a historical precedent of this.
-It is a work by Costa de Beauregard [26, 27] that also provides a retrocausal explanation of quantum nonlocality.
-In it, the outcome of the first performed measurement—of the observable of one of the two space separate subsystems—locally and unitarily goes backward in time until the time the two subsystems were not yet space-separate, where it locally tells the other subsystem the state to assume in view of the future measurement.
-Naturally, we are not dealing with the present time-symmetric form of retrocausality, but the concept is similar.
-Speaking of precedents, of course, the present work is inspired by time-symmetric quantum mechanics [28, 29].
-Furthermore, there have been and there are studies on the possible violations of the causal principle by quantum mechanics.
-However, none of them concerns the quantum computational speedup.
-Their possible relation to the present work, as things are now, would not be evident.
 4 Conclusions
-The present work goes against a long-standing de facto no-retrocausality tradition of physics.
+The present work goes against a long-standing de facto no retrocausality tradition of physics.
 More specifically, since causality could well go backward in time along a unitary quantum evolution that is both objectively and ideally time-reversible, we should say that, in the quantum world, the scientific tradition amounts to asserting the mutual exclusivity between the two possible time directions of causality.
-Let us first address the reliability of the work, a particular issue because of its unconventionality.
-That the classical logic description of the quantum speedup must resort to retrocausality should be obvious.
-Of course, it should be obvious the fact that always locating a ball hidden in one of four drawers by opening a single drawer is logically equivalent to always knowing in advance that the ball is in a pair of drawers.
-To the problem-solver, the only possible source of this information is the solution of the problem she will produce and measure in the future: hence retrocausality.
-We have seen how this immediately generalizes to the full fledged Grover algorithm and a most general form of quantum correlation.
-At this point, we should address the legitimacy of applying classical logic to what the quantum algorithm does. Let us recall our two arguments in favor of this choice.
-The first is that, only involving the measurements of commuting observables (the problem-setting and the solution of the problem), quantum algorithms (and what they do) should be subject to classical logic.
-The other is that all the results obtained under the assumption that quantum algorithms can be described by classical logic are also obtained by time-symmetrizing their ordinary quantum description; this, in turn, is mandatory to satisfy the atemporal symmetry between the simultaneous measurements of two maximally entangled observables implicit in the ordinary quantum description of quantum algorithms.
-We would think that the present retrocausal explanation of the quantum speedup is reliable enough to deserve a discussion despite its destabilizing consequences.
-The flip side of the destabilizing character of the work is its innovative character.
-Let us distinguish between what should be certain and what should be plausible.
-What should be certain: (i) The work provides a unified quantitative explanation of the quantum speedups that is still missing from mainstream quantum mechanics.
-(ii) It provides a theoretical confirmation of the incompleteness of the quantum description of quantum nonlocality sustained by Einstein and others, also providing the local hidden variables envisaged by those authors to complete it.
-(iii) It highlights the teleological character of quantum algorithms; given the negative view that the teleological notion has in today’s science and the scope of this notion, this should be the most important result.
-What should be plausible: It shows that, under a quantum cosmological reading of the Fine-tuned Universe version of the Anthropic Principle, the physical basis of the teleological character of quantum algorithms applies as well to the evolutions of the living.
-In conclusion, the present work should show that there is another possible view of physical reality out there, outside the current view.
-We note that this might not only be a curiosity.
-A revamping of the abandoned notion of the relation between evolution and functional structure of the living might have a profound impact on today’s evolutionary biology and biology.
+Let us first address the reliability of the present unconventional work.
+The logic for which quantum algorithms should be teleological evolutions is simple.
+Always locating a ball hidden in one of four drawers by opening a single drawer, as the quantum algorithm does, is obviously equivalent to knowing in advance that the ball is, in a mutually exclusive way, in one of the three pairs of drawers in which it must be.
+This, of course, implies knowing in advance, in a mutually exclusive way, one of the possible halves of the information about the solution one will produce and measure in the future and using this knowledge to produce the solution with fewer computation steps.
+This attraction from the future is indeed the teleological character of the quantum algorithm.
+It should be completely evident.
+We have seen that this immediately generalizes to any number of drawers and most general forms of quantum correlation.
+The above is also confirmed by the fact that the time-symmetrization of the ordinary quantum description of the quantum algorithms says the same thing and is mandatory to satisfy the atemporal symmetry implicit in the latter description.
+Summarizing, the teleological character of quantum algorithm should be evident, and enough to dispel the belief that teleological evolutions have no physical basis.
 Appendix
@@
 She can then solve the problem in an optimal logical way by opening just one of the two drawers, indeed as in the classical logic description of what the quantum algorithms does.
-Also note that table III diagram goes along with the fact that the unitary part of the quantum algorithm, $\hat{\mathbf{U}}_{1,2}$, takes the opening just one drawer, but does not imply it.
+Also note that table III diagram goes along with the fact that the unitary part of the quantum algorithm, $\hat{\mathbf{U}}_{1,2}$, takes the opening of just one drawer, but does not imply it.
 We know this fact from outside Table III diagram: we know that this is what the unitary part of Grover algorithm takes.